### v1

# Revisiting Cross-Diffusion for Overcrowding Dispersal in Interacting Species System

Abstract
This work introduces and study a novel class of reaction-diffusion systems to model the evolution of two interacting species that disperse to avoid overcrowding. Our approach employs the principle of proximal minimization energy, implemented through a minimum flow proximal algorithm. This framework offers a potential generalization of existing systems commonly used in the theory of segregation for interacting species dynamics. Beyond existence and uniqueness, this approach allows us to capture the dynamic interplay between diffusion rates and concentration gradients of each species, and has the potential to significantly advance our understanding of how cross-diffusion shapes the spatial distribution, coexistence, and pattern formation of multiple species within a system.
1 Introduction and preliminaries
In this study, we focus on a specific type of cross-diffusion system involving overcrowding dispersal of two species. Overcrowding dispersal in interacting species refers to the phenomenon where two or more species, when encountering high population density in a shared environment, exhibit behaviors that encourage them to spread out and reduce their local density. Cross-diffusion systems have gained significant attention in mathematical modeling as they provide a more realistic description of interactions and dynamics between different species in various scientific fields, including ecology, chemistry, and biology. In this context, a cross-diffusion system captures the interplay between multiple species by considering the influence of one species on the diffusion rates of another. It is known that mathematical analysis and numerical simulations of these kind of problems exhibit complex dynamics and patterns which may be handled using cross-diffusion system. The cross-diffusion terms in the system take into account the fact that the diffusion rates of each species are influenced not only by their own concentration gradient but also by the concentration gradient of the other species.
Fo reach $k=1,2,$ let us denote by $\rho_{k}(t,x),$ the concentration of each specie $k,$ at time $t\geq 0$ at position $x\in\Omega$, where $\Omega\subset I\!\!R^{N}$ is a given open bounded domain. Here $N\geq 1$ may be considered equals to $1,2$ or 33 33 for practical situations. A particular example of cross-diffusion system we shall discuss in this paper for the couple of interacting densities $(\rho_{1},\rho_{2})$ is given by the balances law for each $\rho_{k},$ $k=1,2,$
$\partial_{t}\rho_{k}+\nabla\cdot J_{k}[\rho_{1},\rho_{2}]=f_{k}$
where $J_{k}[\rho_{1},\rho_{2}]$ is the mass-flux and $f_{k}$ represents reaction term or simply species supply. To model the dispersal phenomena Gurtin et al. propose in [15] to consider the intuitive constitutive relation for each $J_{k}$ given by
$J_{k}[\rho_{1},\rho_{2}]=\rho_{k}\>\upsilon_{k}[\rho_{1},\rho_{2}],$
relating the mass-flux $J_{k}[\rho_{1},\rho_{2}]$ to the dispersal velocity $\upsilon_{k},$ which measures, among other things, the desire of individuals for each specie $k=1,2,$ to migrate. Moreover, for the dispersal phenomena, they propose to consider, for each $k=1,2,$
$\upsilon_{k}[\rho_{1},\rho_{2}]=-\sigma_{k}\>\nabla(\rho_{1}+\rho_{2}),$
where $0<\sigma_{k}=\sigma_{k}(\rho_{k},\rho_{1}+\rho_{2})$ is assumed to be connected to the dispersivity, so that each specie disperses (locally) toward lower values of total population. Then, the underlying PDE takes the form of cross-diffusion system
(1.1) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\rho_{1}\>\sigma_{1}(\rho_{k},\rho_{1}+\rho_{2})\>\nabla(\rho_{1}+\rho_{2}))=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\rho_{2}\>\sigma_{2}(\rho_{k},\rho_{1}+\rho_{2})\>\nabla(\rho_{1}+\rho_{2}))=f_{2}.\end{array}\right.$
Since the seminal paper by [15], the cross-dffusion system (1.1) was proposed in a series of works with constant dispersivity parameter assumed to be equal to 11 11 ; i.e;
(1.2) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\rho_{1}\>\nabla(\rho_{1}+\rho_{2}))=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\rho_{2}\>\nabla(\rho_{1}+\rho_{2}))=f_{2}.\end{array}\right.$
to study segregation theory (see [7, 8, 10, 16]).
See that, while the system initially appears to exhibit fast-wave propagation (hyperbolic behavior) for each individual density $\rho_{k},$ it surprisingly becomes diffusion-dominated (parabolic behavior) when considering the total cell density $\rho_{1}+\rho_{2}$ in the specific case where both mobilities $\sigma_{1}$ and $\sigma_{2}$ are equal to a single value dependent on the total density $\sigma(\rho_{1}+\rho_{2})$. Interestingly, the segregation property suggests that the system might still retain some diffusion-like behavior even for individual densities.
Our study explores through a new approach how two interacting species, driven to disperse by overcrowding, influence each other’s spatial distribution. This investigation reveals a previously hidden diffusion-like (parabolic) nature of the underlying dynamics, offering valuable new insights into their behavior. Notably, this allows us to address well-posedness questions, such as existence and uniqueness of solutions for more general situations. Moreover, while the hyperbolic formulation presented in equation (1.1) may appear suitable, it is not ideal for scenarios involving the dynamic with Dirichlet boundary conditions, such as crowd motion at an exit. This is because the hyperbolic nature suggests rapid wave propagation, which might not accurately capture the dynamics of controlled outflow through a specific location like an exit. As we will explore further, this presents an intriguing challenge that our approach seeks to address.
We introduce a parabolic version of the dynamic, contrasting with the traditional approaches linked to [15] which models mass fluxes using phenomenological processes. The key difference lies in how we handle the mass flux, $J_{k}[\rho_{1},\rho_{2}]$ for each species $k=1,2.$ Our approach, however, utilizes an optimization framework based on the system’s internal energy, expressed through the total density. Inspired by work on single-species scenarios [19], we employ a steepest descent process to minimize this energy against the diffusion work of each species. This framework allows us to introduce the concept of dispersal potentials, $\eta_{k}.$ These potentials capture the individual’s desire to migrate based on the local densities of both species and the total density, $\rho_{1}+\rho_{2}.$ In other words, we encode the dispersal behavior using specific parameters, $\eta_{k}=\eta_{k}(\rho_{1},\rho_{2},\rho_{1}+\rho_{2}),$ to rewrite the mass flux for each species directly as:
$J_{k}[\rho_{1},\rho_{2}]=-\nabla\eta_{k},\quad\hbox{ in }Q,\hbox{ for each }k=1,2.$
This formulation leads to parabolic equations governing the dynamics of the species densities, $\rho_{k}$:
$\left\{\begin{array}[]{l}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=f_{k},\\ \\ \eta_{k}=\eta_{k}(\rho_{1},\rho_{2},\rho_{1}+\rho_{2})\end{array}\right.\quad\text{in }Q,\text{ for }k=1,2.$
By delving deeper into the optimization process for the system’s internal energy, we uncover a crucial link between the dispersal potentials, $\eta_{k},$ the individual densities, $\rho_{k},$ and the total density, $\rho_{1}+\rho_{2}.$ This connection leads into significant equations, providing valuable tools for analyzing the dynamic behavior of the system. Beyond the standard state equation for nonlinear diffusion phenomena (observed in porous media, Stefan, and Hele-Shaw problems), a unique relationship emerges definitely between the total density and the maximum of the dispersal potentials, $\max(\eta_{1},\eta_{2}).$ Additionally, the dispersal potentials satisfy specific equations for spatial delimitation of region:
(1.3) $\rho_{1}=0\text{ in }[\eta_{1}<\eta_{2}]\quad\text{and}\quad\rho_{2}=0\text{ in }[\eta_{1}>\eta_{2}].$
In particular, combining these equations with the state equation enable the description of dynamics within the drivers regions (see Remark 2:
  • $[\rho_{1}>0\>\&\>\rho_{2}=0]$ and $[\rho_{1}=0\>\&\>\rho_{2}>0]$ : regions exclusively populated by $\rho_{1}$ and $\rho_{2}$, respectively.
  • $[\rho_{1}>0\>\&\>\rho_{2}>0]$: region potentially occupied by both species.
It’s important to note that the regions $[\eta_{1}>\eta_{2}]$ and $[\eta_{2}>\eta_{1}]$ represents the so-called segregated or separated regions, while $[\eta_{1}=\eta_{2}]$ corresponds integrated or desegregated region. This novel approach offers a unique perspective on dispersal dynamics, leveraging an optimization framework to capture the interplay between species and their surrounding environment. It provides a powerful tool for analyzing and understanding the emergence of spatial segregation and the dynamics within these distinct regions.
This framework leverages an internal energy optimization approach instead of direct mass flux modeling. This enables us to effectively connect dispersal potentials, individual species densities, and the total density within the dynamic system. It underpins further analysis in this study. Notably, it enables us to investigate the general scenario where each species dynamics are governed together by their intrinsic overcrowding dispersal forces, external forces acting like transport/convection phenomena and various boundary conditions, including Dirichlet, Neumann, or mixed, and not necessary homogeneous, for each population. To this aim, we introduce the cross-diffusion system
(1.4) $\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k})=f_{k},\quad\hbox{ for }k=1,2,\quad\hbox{ in }Q:=(0,T)\times\Omega,$
where $\eta_{k}=\eta_{k}(\rho_{1},\rho_{2},\rho_{1}+\rho_{2}),$ may be interpreted as dispersal potentials, satisfy conditions (1.3) and need to fill moreover the state equations
(1.5) $\max(\eta_{1},\eta_{2})\in\partial\beta(\rho_{1}+\rho_{2}),$
where $\partial\beta$ represents the sub-differential operator of $\beta,$ which is a given intrinsic application closely connected to the internal energy (see Section 2). Moreover, they enables to handle different boundary conditions for each specie like
(1.6) $\left\{\begin{array}[]{ll}(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Sigma_{N_{k}}:=(0,T)\times\Gamma_{N_{k}}\\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}}:=(0,T)\times\Gamma_{N_{k}}\end{array}\right.$
where $\Gamma_{N_{k}}$ and $\Gamma_{N_{k}}$ constitute partitions of the boundary $\partial\Omega$, and the couples $(g_{k},\pi_{k})$ are given boundary data.
The plan of the paper is the following: in Section 2, we gather some notation and the assumptions and state the main results. We present the existence of a weak solution for the problem (3.37). Then, for the homogeneous case and vanishing transport; i.e. $V\equiv 0,$ we show how one can connect the problem (3.37) to the well established $H^{-1}-$gradient flow dynamic. In Section 3, we show how to establish the model by using minimum flow steepest gradient descent algorithm. In Section 4, we prove the existence of a weak solution by employing steepest gradient descent algorithm and passing to the limit in the approximate solution. Finally, in Section 5, we provide the proof of the connection of the approach with $H^{-1}-$gradient flow in in the homogeneous case.
2 Assumptions and main results
Throughout the paper, we assume that $\Omega$ is bounded regular domain with Lipschitz boundary $\partial\Omega,$ which can be decomposed as follows
$\partial\Omega=\underbrace{\Gamma_{D_{1}}\cup\Gamma_{D_{2}}}_{\Gamma_{D}}\cup\underbrace{\Gamma_{N_{1}}\cup\Gamma_{N_{2}}}_{\Gamma_{N}},$
such that
$\mathcal{H}^{N-1}(\Gamma_{D_{k}})>0\quad\hbox{ and }\quad\overline{\Gamma}_{D_{k}}\cap\overline{\Gamma}_{N_{k}}=\emptyset,\quad\hbox{ for each }k=1,2.$
We consider a function $\beta\>:\>\Omega\times I\!\!R\to I\!\!R$ a carathéodory application satisfying the following assumptions:
  • (H1)
    $\beta(x,0)=0$ a.e. $x\in\Omega$ and, for a.e. $x\in\Omega$, the application
[…]
  • (H2)
    there exists $C,\>M>0$ such that
    $C\>(|r|-M)^{+2}\leq\beta(x,r),\quad\hbox{ for any }r\in I\!\!R.$
Moreover, we fix $\alpha=(\alpha_{1},\alpha_{2})\in I\!\!R^{2}$ and $\sigma=(\sigma_{1},\sigma_{2})\in[L^{\infty}_{+}(\Omega)]^{2}$ such that
(2.7) $0<\min(\alpha_{1},\alpha_{2})\quad\hbox{ and }\quad 0<\min_{x\in\overline{\Omega}}\min(\sigma_{1}(x),\sigma_{2}(x)).$
Our aim is to introduce and study the cross-diffusion system
(2.8) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k})=f_{k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\end{array}\right\}\quad\hbox{ in }Q,$
Here, $\vee$ denotes the maximum operator. $V_{k}$ represents a vector-valued field, and $f_{k}$ is a source term, assumed to be provided for both species. We consider system (2.8) with mixed boundary conditions of the following type:
(2.9) $\left\{\begin{array}[]{ll}(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k})\cdot\nu=\tilde{\pi}_{k}&\hbox{ on }\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}},\end{array}\right.$
where $g_{k}$ and $\tilde{\pi}_{k}$ are given boundary data, for $k=1,2,$ we precise later.
To begin with, for any $1\leq p\leq\infty$ and $\tilde{\Gamma}\subset\partial\Omega$ given, let us remind here some notations and functional sets we use throughout the paper.
    $L^{p}(\Omega)$ denotes the usual Lebesgue spaces endowed with their natural norms. Unless otherwise stated, and without abuse of notation, we will use $\int_{\Omega}h$ to denote the Lebesgue integral over $\Omega,$ i.e.
[…]
    $H^{1}(\Omega)$ denotes the usual Sobolev space endowed with its natural norm
    $\|z\|_{H^{1}(\Omega)}=\left(\int_{\Omega}|z|^{2}+\int_{\Omega}|\nabla z|^{2}\right)^{1/p}.$
    $H^{1}_{\tilde{\Gamma}}(\Omega)$ denotes the closure, in $H^{1}(\Omega)$, of $\mathcal{C}^{1}(\overline{\Omega})$ which are null on $\tilde{\Gamma}.$ In particular, for any $z\in H^{1}_{\tilde{\Gamma}}(\Omega),$ $\gamma(z)=0$ on $\tilde{\Gamma},$ where $\gamma$ denotes the usual trace application defined from $H^{1}(\Omega)\to L^{2}(\partial\Omega).$
    $H^{-1}_{\tilde{\Gamma}}(\Omega)$ denotes the dual space of $H^{1}_{\tilde{\Gamma}}(\Omega).$ We know that (see for instance [19]) , for any $f\in H^{-1}_{\tilde{\Gamma}}(\Omega),$ there exists a couple $(f_{0},\overline{f})\in L^{2}(\Omega)\times L^{2}(\Omega)^{N},$ such that
    (2.10) $\langle f,\xi\rangle_{H^{-1}_{\tilde{\Gamma}}(\Omega),H^{1}_{\tilde{\Gamma}}(\Omega)}=\int_{\Omega}f_{0}\>\xi-\int_{\Omega}\overline{f}\cdot\nabla\xi,\quad\hbox{ for any }\xi\in H^{1}_{\Gamma_{D}}(\Omega).$
    Without abusing, we will use the couple $(f_{0},\overline{f})$ to identify $f\in H^{-1}_{\tilde{\Gamma}}(\Omega)$ and denote the duality bracket by .,.$\langle.,.\rangle_{\Omega}$ ; i.e.
    $\langle f,\xi\rangle_{\Omega}:=\langle f,\xi\rangle_{H^{-1}_{\tilde{\Gamma}}(\Omega),H^{1}_{\tilde{\Gamma}}(\Omega)},\quad\forall\xi\in H^{1}_{\tilde{\Gamma}}(\Omega).$
    $H^{1/2}(\tilde{\Gamma})$ denotes the usual dual Sobolev space given by $\gamma_{/\tilde{\Gamma}}(H^{1}(\Omega)),$ where $\gamma_{/\tilde{\Gamma}}$ is the trace application restricted to $\tilde{\Gamma}.$ We need to define moreover
    $H^{1/2}_{00}(\tilde{\Gamma}):=\Big{\{}\kappa\in L^{2}(\tilde{\Gamma})\>:\>\exists\>\tilde{\kappa}\in H^{1}(\Omega),\>\gamma(\tilde{\kappa})=\kappa\hbox{ on }\tilde{\Gamma}\hbox{ and }\gamma(\tilde{\kappa})=0\hbox{ on }\partial\Omega\setminus\tilde{\Gamma}\Big{\}}.$
    Thanks to the assumptions on $\partial\Omega,$ taking $\tilde{\Gamma}=\Gamma_{D_{k}}$ or $\tilde{\Gamma}=\Gamma_{N_{k}},$ for $k=1,2,$ the space $H^{1/2}_{00}(\tilde{\Gamma})$ coincides with the set of functions belonging to $H^{1/2}(\tilde{\Gamma})$ and vanishing on the boundary $\partial\Omega\setminus\tilde{\Gamma}.$ Remember that this not automatically true for any $\tilde{\Gamma}\subset\partial\Omega$.
    $H^{-1/2}(\tilde{\Gamma})$ denotes the usual the topological dual space of $H^{1/2}_{00}(\tilde{\Gamma}),$ $\left(H^{1/2}_{00}(\tilde{\Gamma})\right)^{\prime}.$ We denote the duality bracket simply by the formal expression $\langle\pi,\kappa\rangle_{\tilde{\Gamma}}$ ; i.e.
    (2.11) $\langle\pi,\kappa\rangle_{\tilde{\Gamma}}:=\langle\pi,\kappa\rangle_{H^{-1/2}(\tilde{\Gamma}),H^{1/2}_{00}(\tilde{\Gamma})},\quad\forall\kappa\in H^{1/2}_{00}(\tilde{\Gamma})\hbox{ and }\pi\in H^{-1/2}(\tilde{\Gamma}).$
    We can define simultaneously the space
    $H_{div}(\Omega):=\Big{\{}\upsilon\in L^{2}(\Omega)^{N}\>:\>\nabla\cdot\upsilon\in L^{2}(\Omega)\Big{\}},$
    where $\nabla\cdot\upsilon$ is taken in $\mathcal{D}^{\prime}(\Omega).$
    For any $\upsilon\in H_{div}(\Omega),$ the normal trace $\upsilon\cdot\nu$ is well defined on $\tilde{\Gamma}$ by duality. More precisely, $\upsilon\cdot\nu\in H^{-1/2}(\tilde{\Gamma})$, and
[…]
    for any $\kappa\in H^{1/2}(\tilde{\Gamma})$ and $\tilde{\kappa}\in H^{1}(\Omega)$ such that $\tilde{\kappa}=\kappa$ on $\tilde{\Gamma},$ and $\tilde{\kappa}=0$ on $\partial\Omega\setminus\tilde{\Gamma}.$
Our first main result concerns existence of weak solution to the cross diffusion system
(2.13) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k}+\overline{f}_{k})=f_{0k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }Q,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k}+\overline{f}_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}},\\ \\ \rho_{k}(0)=\rho_{0k}&\hbox{ in }\Omega,\end{array}\right.$
where
  • $(\pi_{1},\pi_{2})=\pi\in H^{-1/2}(\Gamma_{N_{1}})\times H^{-1/2}(\Gamma_{N_{2}}),$
  • $(g_{1},g_{2})=g\in H^{1/2}_{00}(\Gamma_{D_{1}})\times H^{1/2}_{00}(\Gamma_{D_{2}}),$
  • for each $k=1,2,$ $(f_{0k},\overline{f}_{k})\in L^{2}(Q)\times L^{2}(Q)^{N}$,
  • $(V_{1},V_{2})=V\in L^{\infty}(Q)^{N}\times L^{\infty}(Q)^{N},$
  • $(\rho_{01},\rho_{02})=\rho_{0}\in L^{2}(Q)\times L^{2}(Q).$
Theorem 1
Under the previous assumptions, the system of PDE (2.13) has a solution $(\rho,\eta)$ in the sense that, for each $k=1,2,$ $(\rho_{k},\eta_{k})\in L^{2}_{+}(\Omega)\cap W^{1,2}(0,T;H^{-1}_{\Gamma_{D_{k}}}(\Omega)),$ $\eta_{k}\in L^{2}(0,T;H^{1}(\Omega)),$ $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$
(2.14) $\left\{\begin{array}[]{l}\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0,\end{array}\right.\quad\hbox{ a.e. in }Q,$
(2.15) $\frac{d}{dt}\int_{\Omega}\rho_{k}\>\xi_{k}+\int_{\Omega}(\sigma_{k}\>\nabla\eta_{k}+\rho\>V_{k}+\overline{f}_{k})\cdot\nabla\xi_{k}=\int_{\Omega}f_{0k}\>\xi_{k}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{N_{k}}},\quad\hbox{ in }\mathcal{D}(^{\prime}0,T)$
for any $(\xi_{1},\xi_{2})\in H^{1}_{\Gamma_{D_{1}}}(\Omega)\times H^{1}_{\Gamma_{D_{2}}}(\Omega).$
The key idea behind the connection between $\rho$ and $\eta$ in (2.14) appears through the graph $\partial\mathcal{B}_{\alpha}(x,.),$ where for a.e. $x\in\Omega,$ $\mathcal{B}_{\alpha}(x,.)\>:\>I\!\!R^{2}\to I\!\!R^{+}$ is the application given by
$\mathcal{B}_{\alpha}(x,r)=\left\{\begin{array}[]{ll}\beta(x,\alpha_{1}\>r_{1}+\alpha_{2}\>r_{2})&\hbox{ if }r_{1},\>r_{2}\geq 0\\ \\ +\infty&\hbox{ otherwise}.\end{array}\right.$
Remember that $d\in\partial\mathcal{B}_{\alpha}(x,r)$ if and only if $d,\>r\in I\!\!R^{2}$ and $d\cdot(s-r)\leq\mathcal{B}_{\alpha}(x,s)-\beta(x,r),$ for any $s\in I\!\!R^{2}.$ Working with $\max_{m\in I\!\!R}(m\>q-\beta(x,m))=:\beta^{*}(x,q),$ the so called Legendre transform of $\beta(x,.),$ we have
Proposition 1
For a.e. $x\in\textdiv Omega,$ $\partial\mathcal{B}_{\alpha}(x,.)$ defines a maximal montone graph in $I\!\!R^{2}\times I\!\!R^{2}.$ Moreover, for any $d=(d_{1},d_{2})$ and $r=(r_{1},r_{2})\in I\!\!R^{2}$ we have
$\displaystyle d\in\partial\mathcal{B}_{\alpha}(x,r)\quad$$\Leftrightarrow$ $\displaystyle(r_{1},r_{2})\in\operatorname*{arg\,max}_{s_{1},s_{2}\geq 0}\Big{\{}s_{1}d_{1}+s_{2}d_{2}-\beta(x,\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})\Big{\}}$
[…]
Proof :
It is clear that $\mathcal{B}$ is convex and l.s.c, so that $\partial\mathcal{B}$ is a maximal monotone graph in $I\!\!R^{2}\times I\!\!R^{2}.$ Remember that, $d\in\partial\mathcal{B}_{\alpha}(r)$ is equivalent to say $r_{1},\>r_{2}\>\geq 0$ and
$\displaystyle r_{1}d_{1}+r_{2}d_{2}-\underbrace{\beta(x,\alpha_{1}\>r_{1}+\alpha_{2}\>r_{2})}_{\mathcal{B}_{\alpha}(x,r)}=\underbrace{\max_{s_{1},s_{2}\geq 0}\Big{\{}s_{1}d_{1}+s_{2}d_{2}-\beta(\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})\Big{\}}}_{=:\mathcal{B}_{\alpha}^{*}(x,d)},$
where $\mathcal{B}^{*}$ denotes the Fenchel conjugate of $\mathcal{B}.$ This implies the first equivalence. The remaining equivalence follows readily from the combination of two facts: the inequality
$s_{1}d_{1}+s_{2}d_{2}\leq(\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})\>\frac{d_{1}}{\alpha_{1}}\vee\frac{d_{2}}{\alpha_{2}}\leq\beta(x,\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})+\beta^{*}\left(x,\frac{d_{1}}{\alpha_{1}}\vee\frac{d_{2}}{\alpha_{2}}\right),$
holds to be true for any $d_{1},\>d_{2}\in I\!\!R$ and $s_{1},s_{2}\in I\!\!R^{+},$ and the equality ; i.e. $r_{1}d_{1}+r_{2}d_{2}=\beta(x,\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})+\beta^{*}\left(x,\frac{d_{1}}{\alpha_{1}}\vee\frac{d_{2}}{\alpha_{2}}\right),$ holds to be true for $r_{1},r_{2}\geq 0$ if and only if
(2.17) $r_{1},r_{2}\geq 0,\>r_{1}(d_{1}-d_{1}\vee d_{2})^{+}=r_{2}(d_{2}-d_{1}\vee d_{2})^{+}=0\hbox{ and }\frac{d_{1}}{\alpha_{1}}\vee\frac{d_{2}}{\alpha_{2}}\in\partial\beta\left(x,\alpha_{1}\>r_{1}+\alpha_{2}\>r_{2}\right).$
Thus the results of the proposition $\square$
Remark 1
To illustrate the behavior of the graph $\partial\mathcal{B}_{\alpha}(x,.)$ in $I\!\!R^{2}\times I\!\!R^{2}$ and explore potential relationships between $\rho$ and $\eta,$ we present some explicit formulations of $\partial\mathcal{B}_{\alpha}(x,r)$ and its inverse $(\partial\mathcal{B}_{\alpha}(x,.))^{-1}(d),$ which coincides with $\partial\mathcal{B}_{\alpha}^{*}(x,d)$, for arbitrary $r$ and $d$ in $I\!\!R^{2}$, and a.e. $x\in\Omega.$
  1. 1.
    For a.e. $x\in\Omega,$ $\partial\mathcal{B}_{\alpha}(x,r)$ (resp. $\partial\mathcal{B}^{*}$) may be defined by
    $\partial\mathcal{B}_{\alpha}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}\in\partial\beta(0)\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(d_{1},t)\>:\>d_{1}\in\partial\beta(r_{1}),\>t\leq d_{1}\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,d_{2},)\>:\>d_{2}\in\partial\beta(r_{2}),\>t\leq d_{2}\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{\{}(t,t)\>:\>t\in\partial\beta(r_{1}+r_{2})\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.$
    for any $r\in I\!\!R^{+}\times I\!\!R^{+},$ (resp.
    $\partial\mathcal{B}_{\alpha}^{*}(r)=\left\{\begin{array}[]{ll}\Big{\{}(0,t)\>:\>t\in\partial\beta^{*}(d_{2})\Big{\}}&\hbox{ if }d_{1}<d_{2}\\ \Big{\{}(t,0)\>:\>t\in\partial\beta^{*}(d_{1})\Big{\}}&\hbox{ if }d_{1}>d_{2}\\ \Big{\{}(t_{1},t_{2})\>:\>t_{1}\geq 0,\>t_{2}\geq 0,\>t_{1}+t_{2}\in\partial\beta^{*}(d_{1})\Big{\}}&\hbox{ if }d_{1}=d_{2}\\ \end{array}\right.$
    for any $\alpha\in I\!\!R\times I\!\!R$).
  2. 2.
    In the case where $\beta$ (resp. $\beta^{*}$) is differentiable, for a.e. $x\in\Omega,$ $\partial\mathcal{B}_{\alpha}(x,r)$ (resp. $\partial\beta^{*}(x,d)$) may be defined by
    $\partial\mathcal{B}_{\alpha}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(\beta^{\prime}(r_{1}),t)\>:\>t\leq\beta^{\prime}(r_{1})\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,\beta^{\prime}(r_{2}))\>:\>t\leq\beta^{\prime}(r_{2})\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{(}\beta^{\prime}(r_{1}+r_{2}),\beta^{\prime}(r_{1}+r_{2})\Big{)}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.$
    for any $r\in I\!\!R^{+}\times I\!\!R^{+}$ (resp.
    $\partial\mathcal{B}_{\alpha}^{*}(d)=\left\{\begin{array}[]{ll}(0,(\beta^{*})^{\prime}(d_{2})&\hbox{ if }d_{1}<d_{2}\\ ((\beta^{*})^{\prime}(d_{1}),0)&\hbox{ if }d_{1}>d_{2}\\ \Big{\{}(t_{1},t_{2})\>:\>t_{1}\geq 0,\>t_{2}\geq 0,\>t_{1}+t_{2}=(\beta^{*})^{\prime}(d_{1})&\hbox{ if }d_{1}=d_{2}\\ \end{array}\right.$
    for any $d\in I\!\!R\times I\!\!R$).
  3. 3.
    For instance in the case where $\beta(r)=\frac{1}{m+1}r^{m+1}$, the porous medium framework, $\partial\mathcal{B}_{\alpha}(x,r)$ may be defined by
    $\partial\mathcal{B}_{\alpha}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(r_{1}^{m},t)\>:\>t\leq r_{1}^{m}\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and}r_{2}=0\\ \Big{\{}(t,r_{2}^{m})\>:\>t\leq r_{2}^{m}\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and}r_{2}>0\\ \Big{(}(r_{1}+r_{2})^{m},(r_{1}+r_{2})^{m}\Big{)}&\hbox{ if }r_{1}>0\hbox{ and}r_{2}>0,\end{array}\right.$
    for any $r\in I\!\!R^{+}\times I\!\!R^{+}$
  4. 4.
    In the case where $\beta(r)=I\!\!I_{[0,1]}$, which corresponds to crowed motion models, $\partial\mathcal{B}_{\alpha}(x,r)$ may be defined by
    $\partial\mathcal{B}_{\alpha}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(\eta_{1},t)\>:\>t\leq\eta_{1}^{m},\>(1-r_{1})\>\eta_{1}=0\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,\eta_{2})\>:\>t\leq\eta_{2},\>(1-r_{2})\>\eta_{2}=0\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{\{}(t,t)\>:\>\>(1-r_{1}-r_{2})\>t=0\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.$
    for any $r\in I\!\!R^{+}\times I\!\!R^{+}$
Remark 2
Thanks to Remarks, the parabolic nature of the dynamics in the system (2.13) becomes clear. Indeed, assume for instance that $\overline{f}$ is equal to $0,$ and $\beta$ is differentiable. On sees that the dynamics formally split into three driver regions, each exhibiting specific behavior (assuming they are sufficiently regular):
$\left\{\begin{array}[]{lll}\rho_{1}>0,\>\rho_{2}=0\\ \\ \partial_{t}\rho_{1}-\nabla\cdot(\alpha_{1}\>\sigma_{1}\>\nabla\beta^{\prime}(\alpha_{1}\>\rho_{1})+\rho_{1}V_{1})=f_{1}\\ -\nabla\cdot(\sigma_{2}\>\nabla\eta_{2})=f_{2},\>\frac{\eta_{2}}{\alpha_{2}}\leq\frac{1}{\alpha_{1}}\beta^{\prime}(\alpha_{1}\>\rho_{1})\end{array}\right.$
$\left\{\begin{array}[]{lll}\rho_{1}=0,\>\rho_{2}>0\\ \\ -\nabla\cdot(\sigma_{1}\>\nabla\eta_{1})=f_{1},\>\frac{\eta_{1}}{\alpha_{1}}\leq\frac{1}{\alpha_{2}}\beta^{\prime}(\alpha_{2}\>\rho_{2})\\ \partial_{t}\rho_{2}-\nabla\cdot(\alpha_{2}\>\sigma_{2}\>\nabla\beta^{\prime}(\alpha_{2}\>\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
$\left\{\begin{array}[]{lll}\partial_{t}(\rho_{1}-\rho_{2})+\nabla\cdot(\rho_{1}\>V_{1}-\rho_{2}\>V_{2})=f_{1}-f_{2}\\ \\ \partial_{t}\rho_{1}-\nabla\cdot(\alpha_{1}\>\sigma_{1}\>\nabla\beta^{\prime}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})+\rho_{1}V_{1})=f_{1}\\ \partial_{t}\rho_{2}-\nabla\cdot(\alpha_{2}\>\sigma_{2}\>\nabla\beta^{\prime}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
The system in equation (2.13) exhibits similarities to the commonly used system in the literature shown in
(2.18) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\alpha_{1}\>\sigma_{1}\>\rho_{1}\>\nabla\varphi^{\prime}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})+\rho_{1}\>V_{1})=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\alpha_{2}\>\sigma_{2}\>\rho_{2}\>\nabla\varphi^{\prime}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})+\rho_{1}\>V_{1})=f_{2}.\end{array}\right.$
Both systems share a similar behavior in the segregated regions, $S_{1},$ and $S_{2}$ as long as the function $\varphi$ satisfies $\beta(r)=r\>\varphi(r)-r,$ for any $r\in I\!\!R.$ However, there are key differences. For instance in the case where $\alpha_{k}=1,$ for $k=1,2,$ :
- Compatibility with source/reaction terms: Equation (2.13) appears to be more compatible and adaptable to include more possibilities for source/reaction terms compared to (2.18), even in segregated regions.
- Parabolic nature: In the aggregated region, $S,$ the systems diverge significantly. While (2.13) exhibits a fully parabolic regime with respect to all three variables $\rho_{1},$ $\rho_{2}$ and $\rho_{1}+\rho_{2},$ equation (2.18) is only parabolic with respect $\rho_{1}+\rho_{2}.$ Indeed, a formal analysis on the behavior of the solution of (2.18) into the three driver regions (assuming again they are sufficiently regular) implies that :
$\left\{\begin{array}[]{lll}\rho_{1}>0,\>\rho_{2}=f_{2}=0\\ \\ \partial_{t}\rho_{1}-\nabla\cdot(\sigma_{1}\>\rho_{1}\>\nabla\alpha^{\prime}(\rho_{1})+\rho_{1}V_{1})=f_{1}\end{array}\right.$
$\left\{\begin{array}[]{lll}\rho_{1}=f_{1}=0,\>\rho_{2}>0\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\sigma_{2}\>\rho_{2}\nabla\alpha^{\prime}(\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
$\left\{\begin{array}[]{lll}\partial_{t}\rho_{1}-\nabla\cdot(\sigma_{1}\>\rho_{1}\nabla\alpha^{\prime}(\rho_{1}+\rho_{2})+\rho_{1}V_{1})=f_{1}\\ \partial_{t}\rho_{2}-\nabla\cdot(\sigma_{2}\>\rho_{2}\nabla\alpha^{\prime}(\rho_{1}+\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
Remark 3
Without any external factors influencing the system ($f\equiv 0$ and $V\equiv 0$), in the case where $\sigma_{k}=\alpha_{k}=1,$ for $k=1,2,$ we expect the solution of (2.13) to follow this rule
$\left\{\begin{array}[]{lll}\rho_{1}>0,\>\rho_{2}=0\\ \\ \partial_{t}\rho_{1}-\Delta\beta^{\prime}(\rho_{1})=0\\ -\Delta\eta_{2}=0,\>\eta_{2}\leq\beta^{\prime}(\rho_{1})\end{array}\right.$
$\left\{\begin{array}[]{lll}\rho_{1}=0,\>\rho_{2}>0\\ \\ -\Delta\nabla\eta_{1}=0,\>\eta_{1}\leq\beta^{\prime}(\rho_{2})\\ \partial_{t}\rho_{2}-\Delta\beta^{\prime}(\rho_{2})=0\end{array}\right.$
$\left\{\begin{array}[]{lll}\rho_{1}-\rho_{2}=\rho_{01}-\rho_{02}\\ \\ \partial_{t}\rho_{1}-\Delta\beta^{\prime}(\rho_{1}+\rho_{2})=0\\ \partial_{t}\rho_{2}-\Delta\beta^{\prime}(\rho_{1}+\rho_{2})=0\end{array}\right.$
Particularly for homogeneous boundary conditions, we expect $\eta_{1}=0$ (respectively, $\eta_{2}=0$) in region $S_{2}$ (respectively, $S_{1}$). Additionally, given that $\rho_{1}-\rho_{2}=\rho_{01}-\rho_{02}$ in region $S,$ we anticipate $\rho_{1}=\rho_{2}$ outside the combined support of the initial densities $\rho_{01}$ and $\rho_{02}.$
An interesting future direction of this approach lies in establishing a formal connection between the dynamics in equation (2.13) and $H^{-1}-$gradient flow, mirroring the relationship observed in single-species models. We demonstrate this connection in the context of two species, thereby providing a general abstract framework for cross-diffusion systems arising from overcrowding dispersal phenomena. For each $k=1,2,$ we consider the dual Sobolev space $H^{-1}_{\Gamma_{D_{k}}}(\Omega)$. It is not difficult to see that
(2.19) $\|f\|_{H^{-1}_{\Gamma_{D_{k}}}}=\min_{\omega\in L^{2}(\Omega)^{N}}\left\{\left(\int_{\Omega}\sigma_{k}\>|\omega|^{2}\right)^{1/2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega+\overline{f}_{k})=f_{0k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega+\overline{f}_{k})\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\},$
for any $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ defines a norm on $H^{-1}_{\Gamma_{D_{k}}}(\Omega).$ This norm is clearly associated with the inner product
$\langle f,g\rangle_{k,\Omega}:=\int_{\Omega}\sigma_{k}\>\phi^{f}\cdot\phi^{g},\quad\hbox{ for any }f,g\>\in H^{-1}_{\Gamma_{D_{k}}}(\Omega)$
where $\phi^{f}$ and $\phi^{g}$ are given by
$\phi^{f}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}|\sigma_{k}\>\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega+\overline{f}_{k})=f_{0k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega+\overline{f}_{k})\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\}$
and
$\phi^{g}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}\sigma_{k}\>|\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega+\overline{g}_{k})=g_{0k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega+\overline{g}_{k})\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\}.$
Remember here (see for instance Proposition 2), that there exists a unique $\eta^{f}_{k}\in H^{1}_{\Gamma_{D_{k}}}\!\!(\Omega)$ (resp. $\eta^{g}_{k}\in H^{1}_{\Gamma_{D_{k}}}\!\!(\Omega)$) such that $\phi^{f}_{k}=\nabla\eta^{f}_{k}$ ( resp. $\phi^{g}_{k}=\nabla\eta^{g}_{k}$) a.e. in $\Omega.$ This implies that $H^{-1}_{\Gamma_{D_{k}}}\!\!(\Omega)$ equipped with the norm .$\|.\|_{H^{-1}_{\Gamma_{D_{k}}}}$ and the inner product .,.$\langle.,.\rangle_{k,\Omega}$ is a Hilbert space.
Now let us consider the space $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega),$ endowed with the norm
$\|f\|_{H^{-1}_{\Gamma_{D}}}=\left(\|f_{1}\|_{H^{-1}_{\Gamma_{D_{1}}}}^{2}+\|f_{2}\|_{H^{-1}_{\Gamma_{D_{2}}}}^{2}\right)^{1/2},\quad\hbox{ for any }f=(f_{1},f_{2})\in H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega).$
This is in turn a Hilbert space, and the the associate inner product is given by
$[f,g]_{\Omega}=\langle f_{1},g_{1}\rangle_{1,\Omega}+\langle f_{2},g_{2}\rangle_{k,\Omega},\quad\hbox{ for any }f,g\in H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega).$
Now, let us define the functional $\mathcal{E}\>:\>H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)\to[0,\infty]$ by
(2.20) $\mathcal{E}(\rho):=\left\{\begin{array}[]{ll}\int_{\Omega}\beta(.,\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})&\hbox{ if }\rho_{1},\>\rho_{2},\>\beta(.,\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})\in L^{1}_{+}(\Omega)\\ \\ +\infty&\hbox{ otherwise }.\end{array}\right.$
Thanks to assumptions $(H1)$ and $(H2)$, one sees that $\mathcal{E}$ is convex and l.s.c. in $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega).$ This result implies that the sub-differential of $,\displaystyle\mathcal{E},$$\partial\mathcal{E},$ is well-defined and forms a maximal monotone graph within $\left(H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega),\|-\|_{H^{-1}_{\Gamma_{D}}}\right).$ Thus we can consider the $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow problem
(2.21) $\left\{\begin{array}[]{ll}\frac{d}{dt}\rho(t)+\partial\mathcal{E}(\rho(t))\ni f(t)&\hbox{ a.e. }t\in(0,T)\\ \\ \rho(0)=\rho_{0}.\end{array}\right.$
Our aim now, is to establish the connection the problems (2.13) and (2.21). This is summarized in the following theorem.
Theorem 2
Under the assumptions $g\equiv 0$, $\pi\equiv 0$, $V\equiv 0,$ $f\in BV\left(0,T;H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)\right)$, let us consider $(\rho,\eta)$ be a weak solution of the system of PDE (2.13). Then, $\rho$ coincides with the unique strong solution of (2.21).
Remark 4
  1. 1.
    Theorem 2 bridges a connection between the cross-diffusion system (3.37) and the well-established theory of gradient flows in Hilbert spaces. Furthermore, the specific form of the internal energy (2.20) allows us to definitively connect the differential operator $\mathcal{A}$ (defined precisely in Proposition 5) governing (3.37) to the dispersal dynamics of overcrowding for two species. As to the connection with systems of type (1.2) and (2.18) commonly use din the literature, we de believe that is likely more nuanced and may hold true in specific cases, particularly when solutions exist, such as in fully segregated scenarios.
  2. 2.
    Theorem 2 also yields other crucial results which uses the equivalence between weak solutions and $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow solutions. We focus here on the uniqueness and large time behavior which could be of great importance for the applications :
    1. (a)
      Uniqueness : Since, in the case where $V\equiv 0,$ a weak solution to equation (3.37) corresponds to the strong solution of (2.21), we deduce that they are unique. Notably, this allows to infer the uniqueness of weak solutions (whenever they exist) in fully segregated scenarios.
    2. (b)
      Asymptotic behavior $t\to\infty$ : thanks to the theory of gradient flows in Hilbert spaces, we see that if $f\in L^{2}(0,\infty;H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)),$ satisfies
      $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-\lim_{t\to\infty}f(t)=f_{\infty},$
      then the solution $\rho(t)$ converges uniformly in $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)$ to $\rho,$ as $t\to\infty,$ where $\rho$ is such that there exists $\eta\in[H^{1}(\Omega)]^{2}$ for which the couple $(\rho,\eta)$ is a solution of the stationary problem
      (2.22) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{\infty k})=f_{\infty 0k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }Q,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{\infty k})\cdot\nu=\pi_{k}&\hbox{ on}\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}}.\end{array}\right.$
  3. 3.
    In the case where $V\not\equiv 0,$ the $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow theory is no longer applicable. The uniqueness for weak solutions of the problem (3.37) remains to be a challenging open problem. Remember for the case of one specie the uniqueness may follows through contraction principles in $L^{1}$ and/or in $\mathcal{W}_{2}-$Wasserstein space (one can see [18, 22] for discussions and references in this direction). It is not clear yet for us how to deal with $L^{1}-$theory for cross-diffusion of the type (3.37).
  4. 4.
    We focus here in the case of source term $f=f(t,x).$ However, one sees that one can use the results of Theorem 2 also for treating more general reaction terms $f=f(t,x,\rho,\eta).$
  5. 5.
    At first glance, at least for overcrowding dynamics of one specie, the theories of gradient flows in $H^{-1}$ and $\mathcal{W}_{2}-$Wasserstein space appear to be similar, providing comparable dynamic behaviors through more or less similar internal energies. It is surprising to note here that, this seems not to be the case for two species. Indeed, recent works suggests that the theory of gradient flows in $\mathcal{W}_{2}-$Wasserstein space appears to be closer to the model (2.18) à la Gurtin et al. (cf. [15]) than the $H^{-1}$-theory, which provides complete and general results for (3.37) and seems to generalize (2.18). We believe that further developments using $\mathcal{W}_{2}$ and $H^{-1}$ frameworks should be carried out to complete or explain more rigorously the links between both approaches in the case of two species.
3 Modeling through proximal minimization algorithm
3.1 Preliminaries
As introduced earlier, our approach to cross-diffusion modeling relies on a proximal minimization algorithm. This algorithm, commonly used in the context of nonlinear semigroups within Banach/Hilbert spaces, emerges from the Euler-implicit scheme. It also plays a crucial role in the JKO scheme, minimizing movement within the gradient flow in Wasserstein space, and showcases connections to the gradient descent method used for optimization. In a previous work [19], we prove its application to establish a general nonlinear diffusion equation for a single species. This paper aims to extend this approach to the case of two interacting species. We refer readers to [19] for further details, discussions, development, and references related to this method.
Recall that proximal minimization algorithms target minimizing the system’s internal energy against the work associated with transitions from a given state to a next one. Following [19], we propose employing the minimum flow process to measure this work. This process enables to entail the appropriate dynamic procedure that continuously updates the system’s state, aiming to decrease its internal energy through a nonlinear diffusion process.
More formally, let $X$ be a state space and $\mathcal{E}(\rho)$ be the internal energy of a given system at state $\rho\in X.$ At each time $t>0,$ the dynamic system tends simultaneously to decrease its internal energy and to minimize the work required to move from state $\rho(t)$ to state $\rho(t+h).$ This can be achieved by solving the following optimization problem (proximal energy) at each time step:
$\min_{\rho\in X}\Big{\{}\mathcal{E}(\rho)+W^{h}(\rho(t),\rho)\Big{\}},$
where $W^{h}(\rho(t),\rho)$ is a measure of the work required to move from state $\rho(t)$ to state $\rho,$ throughout small time step $h>0.$ The solution to this optimization problem $\rho(t+h)$ is the next state of the system. To the point, one can consider the $\tau-$time steps defined by $t_{0}=0<t_{1}<\cdots<t_{n}<T$, and the sequence of piecewise constant curves
$\rho_{\tau}=\sum_{i=1}^{n}\rho_{i}\>\chi_{]t_{i-1},t_{i}]},$
where
(3.23) $\rho_{i}=\hbox{argmin}_{\rho\in X}\Big{\{}\mathcal{E}(\rho)+W^{\tau}(\rho_{i-1},\rho)\Big{\}},\quad\hbox{ for }i=1,...n,\hbox{ with }\rho_{0}=\rho(0).$
Then, we expect the limit as $\tau\to 0$ of $\rho_{\tau}$ converges to the solution of a continuous evolution problem. Working within the framework of Hilbert space $(X,\|-\|),$ the general theory suggests to connect $W^{\tau}(\rho(t),\rho)$ to $\|\rho(t)-\rho\|^{2}/2\tau$ to obtain a representation of the solution through the gradient flow equation:
(3.24) $\left\{\begin{array}[]{ll}\rho_{t}(t)+\partial\mathcal{E}(\rho(t))\ni 0\quad\hbox{ in }(0,T)\\ \\ \rho(0)\hbox{ is given},\end{array}\right.$
where $\partial\mathcal{E}(\rho)$ denote sthe sub-differential (possibly multi-valued map) of the functional $\rho\in X\to\mathcal{E}(\rho)\in(-\infty,\infty].$ The expression (3.23) can be related to both to the resolvent associated with the operator $\partial\mathcal{E}$ and Euler-Implicit discretisation of (3.24). The evolution problem in turn is a gradient flow in a Hilbert space which describes the evolution of a continuous curve $t\to\rho(t)\in X$ that follows the direction of steepest descent of the functional $\mathcal{E}$ (cf. [9]). It is possible also to consider a more general situation where $(X,d)$ is metric space and $W^{h}$ is build on the distance $d$. In this case, the expression (3.23) reveals the minimizing movement scheme à la De Giorgi (see [4]). The approach resulted in a broad theory of gradient flows in a metric space where the derivatives $\rho_{t}$ and $\partial\mathcal{E}(\rho)$ need to be interpreted in an appropriate way in $(X,d)$ (cf. the book [4]).
For instance, the porous medium-like equation
(3.25) $\partial_{t}\rho-\Delta\rho^{m}=0$
in a bounded domain arises in various applications, especially in biology, to represent the evolution of species as they strive to minimize their internal energy $\mathcal{E}(\rho),$ which depends inherently on the density $\rho.$ Since the works [26] and [20], it has been established that the PDE (3.25) subject to homogeneous Neuman boundary condition can be derived by employing $\mathcal{E}(\rho)=\frac{1}{m-1}\int_{\Omega}\rho^{m}$ in the context of $W_{2}-$Wasserstein distance by setting $W^{\tau}(\rho,\tilde{\rho})=\frac{1}{2\tau}\mathcal{W}_{2}(\rho,\tilde{\rho})^{2}$. It is well known that it can be also derived by utilizing the steepest descent algorithm for the internal energy $\mathcal{E}(\rho)=\frac{1}{m+1}\int_{\Omega}\rho^{m+1}$ in the context of $H^{-1}-$norm. This is achieved by setting $W^{\tau}(\rho,\tilde{\rho})=\frac{1}{2\tau}\|\rho-\tilde{\rho}\|_{H^{-1}}^{2}$ (cf. [6, 19]). These concepts can be expanded to accommodate some monotone nonlinearity $\eta(r)$ instead of $r^{m},$ both for $W_{2}-$Wasserstein distance and also $H^{-1}-$norm cf. [22, 11, 12, 21]) frameworks. These frameworks can be also extended to nonlinear diffusion operators (see [1] for the $W_{c}-$Wasserstein distance framework, and the recent work [19] for the the case a generalized dual Sobolev approach).
While similar at first glance, analyzing the dynamics of two interacting species using gradient flow becomes challenging. This paper focuses on a system with two species $(k=1,2)$, represented by densities $\rho_{1},$ $\rho_{2},$ where the state space is $X=X_{1}\times X_{2}.$ To capture the overcrowding dispersal dynamic, we consider an internal energy function dependent on the sum of both densities ; i.e. $\mathcal{E}(\rho)=\tilde{\mathcal{E}}(\rho_{1}+\rho_{2}).$ We employ a proximal optimization algorithm to construct a sequence of density pairs $(\rho_{1},\rho_{2})$ starting from initial densities $(\rho_{01},\rho_{02})$. The algorithm iteratively updates the densities using the following steepest descent formula:
(3.26) $\rho_{i}=\hbox{argmin}_{\rho\in X}\Big{\{}\tilde{\mathcal{E}}(\rho_{1}+\rho_{2})+W^{\tau}_{1}(\rho_{1}^{i-1},\rho_{1})+W^{\tau}_{2}(\rho_{2}^{i-1},\rho_{2})\Big{\}}).$
Here, $W^{\tau}_{k}$ measures the work required for each species $k$ to move from state $\rho_{k}^{i-1}$ to state $\rho_{k},$ in a small time step $\tau>0.$ For PME nonlinear diffusion process, previous works (see [10, 22, 13]) suggest that when $W^{\tau}_{k}$ is defined by the $\mathcal{W}_{2}$-Wasserstein distance and $\mathcal{E}(\rho)=\frac{1}{m-1}\int_{\Omega}(\rho_{1}+\rho_{2})^{m}$, this approach leads to a cross-diffusion system:
(3.27) $\left\{\begin{array}[]{ll}{\rho_{1}}_{t}-\frac{m}{m-1}\>\nabla\cdot\left(\rho_{1}\>\nabla(\rho_{1}+\rho_{2})^{m}\right)=0\\ \\ {\rho_{2}}_{t}-\frac{m}{m-1}\>\nabla\cdot\left(\rho_{2}\>\nabla(\rho_{1}+\rho_{2})^{m}\right)=0\end{array}\right.\quad\hbox{ in }Q.$
This system, subject to homogeneous Neumann boundary conditions and initial data, describes how both species avoid overcrowding regions through hyperbolic dynamics like in (1.2). However, in this paper we deviate from the previous approach and proposes using the minimum flow problem to measure the work required for state transitions. We consider:
(3.28) $\rho_{i}=\hbox{argmin}_{\rho\in X}\Big{\{}\tilde{\mathcal{E}}(\rho_{1}+\rho_{2})+\frac{1}{2\tau}\>I_{1}(\rho_{1}^{i-1},\rho_{1})+\frac{1}{2\tau}\>I_{2}(\rho_{2}^{i-1},\rho_{2})\Big{\}},$
where $I_{k}(\rho_{k}^{i-1},\rho_{k})$ minimizes the following cost function:
(3.29) $\frac{1}{2}\int_{\Omega}\sigma_{k}\>|\omega|^{2}-\langle\omega\cdot\nu,g_{k}\rangle_{\Gamma_{D_{k}}}$
over $\omega\in L^{2}(\Omega)^{N}$ subject to the constraint:
$-\nabla\cdot(\sigma_{k}\>\omega)=\rho_{k}^{i-1}-\rho_{k},\quad\text{ in }\Omega,$
with potential boundary conditions for $\sigma_{k}\>\omega\cdot\nu$ on $\partial\Omega\setminus\Gamma_{N_{k}}.$ The goal of $I_{k}(\rho_{k}^{i-1},\rho_{k})$ is to find the most efficient ”traffic scheme” for transferring mass from $\rho_{k}^{i-1}$ to $\rho_{k}.$ This scheme considers both the quadratic work of the flow $\omega$ between densities $\rho_{k}^{i-1}$ and $\rho_{k},$ and the charge $g_{k}$ for $\phi\cdot\nu$ at at specific boundary locations $\Gamma_{D}.$ Since the balance equation, (3.29) captures the essence of mass transfer, the transfer fee effectively models the diffusion process with boundary conditions, enabling analysis of overcrowding dynamics for two interacting species. Interestingly, unlike the single-species case, we’ll see that taking for instance $\mathcal{E}(\rho)=\frac{1}{m+1}\int_{\Omega}(\rho_{1}+\rho_{2})^{m+1}$ this approach leads to a fundamentally different cross-diffusion system generalizing in some sense (3.27) (see Remark 2). This new system is parabolic in nature :
(3.30) $\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=0,\quad\hbox{ for}k=1,2,\quad\hbox{ in }Q:=(0,T)\times\Omega,$
where the potential $\rho_{k}$ and $\eta_{k}$ satisfies the spatial delimitation of region conditions (1.3) and the state equation
$\eta_{1}\vee\eta_{2}=(\rho_{1}+\rho_{2})^{m}.$
Moreover, the approach incorporates non-homogeneous boundary conditions like those in equation (1.6). In the following, section we precise the assumptions and give the rigorous proofs for the results in a general case. Then we show how one can use this approach to prove the results of Section 2. We show also how to connect the approach with gradient flow in dual Sobolev space in the homogeneous case.
3.2 Notations
For the sake of simplicity and ease of presentation, we adopt the following formal notations:
    $L^{2}_{+}(\Omega)$ denotes the set on nonegative Lebesgue integrable functions of $L^{2}(\Omega).$
    $[H^{1}_{\Gamma_{D}}(\Omega)]^{2}:=H^{1}_{\Gamma_{D_{1}}}(\Omega)\times H^{1}_{\Gamma_{D_{2}}}(\Omega).$
    $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}=H^{-1}_{\Gamma_{D_{1}}}(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}(\Omega)$
    $[H^{1/2}(\Gamma_{D})]^{2}:=H^{1/2}(\Gamma_{D_{1}})\times H^{1/2}(\Gamma_{D_{2}})$ And $[H^{1/2}_{00}(\Gamma_{D})]^{2}:=H^{1/2}_{00}(\Gamma_{D_{1}})\times H^{1/2}_{00}(\Gamma_{D_{2}}).$ Moreover, for any $\xi\in[H^{1}(\Omega)]^{2}$ and $g\in[H^{1/2}(\Gamma_{D})]^{2},$ we say $\xi_{/\Gamma_{D}}=g,$ if and only if
    $\xi_{1/\Gamma_{D_{1}}}=g_{1}\hbox{ and }\xi_{2/\Gamma_{D_{2}}}=g_{2}.$
    $[H^{-1/2}(\Gamma_{D})]^{2}:=H^{-1/2}(\Gamma_{D_{1}})\times H^{-1/2}(\Gamma_{D_{2}}).$ Moreover, for any $h=(h_{1},h_{2})\in[H^{-1/2}(\Gamma_{D_{1}})]^{2}\times H^{-1/2}(\Gamma_{D_{2}})$ and $g=(g_{1},g_{2})\in[H^{1/2}_{00}(\Gamma_{D})]^{2}$ we use the notation
    $[h,g]_{\Gamma_{D}}\>:=\langle h_{1},g_{1}\rangle_{{H^{-1/2}(\Gamma_{D_{1}}),H^{1/2}(\Gamma_{D_{1}})}}+\langle h_{2},g_{2}\rangle_{{H^{-1/2}(\Gamma_{D_{2}}),H^{1/2}(\Gamma_{D_{2}})}}.$
    Without abusing, for any $\phi=(\phi_{1},\phi_{2})\in H_{div}(\Omega)\times H_{div}(\Omega),$ we’ll use again the notation $[\phi\cdot\nu,g]_{\Gamma_{D}}$ to point out the sum of the duality products $<\phi_{k}\cdot\nu,g_{k}>$ on $\Gamma_{D_{k}}$ ; i.e.
    $\begin{array}[]{l}[\phi\cdot\nu,g]_{\Gamma_{D}}:=\langle\phi_{1}\cdot\nu,g_{1}\rangle_{{H^{-1/2}(\Gamma_{D_{1}}),H^{1/2}(\Gamma_{D_{1}})}}+\langle\phi_{2}\cdot\nu,g_{2}\rangle_{{H^{-1/2}(\Gamma_{D_{2}}),H^{1/2}(\Gamma_{D_{2}})}}.\end{array}$
At last, for any $\phi=(\phi_{1},\phi_{2})\in L^{2}(\Omega)^{N}\times L^{2}(\Omega)^{N}$ and $\sigma=(\sigma_{1},\sigma_{2})\in[L^{\infty}(\Omega)]^{2},$ we use the notations
$\sigma\>\phi=(\sigma_{1}\phi_{1},\sigma_{2}\>\phi_{2})$
$[\nabla\cdot\sigma\phi]:=(\nabla\cdot(\sigma_{1}\>\phi_{1}),\nabla\cdot(\sigma_{2}\>\phi_{2})),\quad\forall\>\phi=(\phi_{1},\phi_{2})\in L^{2}(\Omega)^{N}\times L^{2}(\Omega)^{N}$
and
$[\sigma\>\nabla\eta]:=(\sigma_{1}\>\nabla\eta_{1},\sigma_{2}\>\nabla\eta_{2}),\quad\forall\eta=(\eta_{1},\eta_{2})\in[H^{1}(\Omega)]^{2}.$
Then, we use the notations $F_{\sigma}[\phi],$ for the quantity
(3.31) $F_{\sigma}[\phi]=\frac{1}{2}\left(\sigma_{1}\>|\phi_{1}|^{2}+\sigma_{2}|\phi_{2}|^{2}\right),\quad\hbox{ for any }\phi\in[L^{2}(\Omega)^{N}]^{2}$
and $\beta[\rho]$ for
$\beta_{\alpha}[\rho]:=\beta(.,\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2}),\quad\hbox{ for any }\rho\in[L^{2}(\Omega)]^{2}.$
3.3 Proximal minimization for cross-diffusion modeling
Let
(3.32) $\pi=(\pi_{1},\pi_{2})\in[H^{-1/2}(\Gamma_{N})]^{2}\quad\hbox{ and }\quad g=(g_{1},g_{2})\in[H^{1/2}_{00}(\Gamma_{D})]^{2},$
be given. We denote by $\tilde{g}$ the $[H^{1}(\Omega)]^{2}$ function such that $\tilde{g}=0$ on $\Gamma_{N}$ and $\tilde{g}=g$ on $\Gamma_{D}.$
To define the transition work rigorously, for any $\mu\in[L^{2}(\Omega)]^{2}$ and $\chi\in[L^{2}(\Omega)^{N}]^{2}$, we consider
$\mathcal{A}_{\pi}^{\chi}[\mu]_{k}=\Big{\{}\omega\in L^{2}(\Omega)^{N},\>-\nabla\cdot(\sigma_{k}\>\omega)=\mu_{k}+\nabla\cdot\chi_{k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\omega+\chi_{k})\cdot\nu=\pi_{k}\hbox{ on }\Gamma_{N_{k}}\Big{\}},\quad k=1,2.$
That is, for any $k=1,2,$ $\omega\in\mathcal{A}_{\pi}^{\chi}[\mu]_{k}$ if and only if $\omega\in L^{2}(\Omega)^{N}$ and
[…]
In some sense, for each $k=1,2,$ $\mathcal{A}_{\pi}^{\chi}[\mu]_{k}$ is the set of all mass fluxes which balance the mass $\mu_{k},$ under the action of the external force $\chi_{k}$ and the boundary action $\pi_{k}$ on $\Gamma_{k}.$ Recall that (see for instance [19]) that, for any $k=1,2,$ we have
$\mathcal{A}_{\pi}^{\chi}[\mu]_{k}\neq\emptyset,\quad\hbox{ for any }\mu\in[L^{2}(\Omega)]^{2}\hbox{ and }\chi\in[L^{2}(\Omega)^{N}]^{2}.$
Now, for any $\mu=(\mu_{1},\mu_{2})\in[L^{2}(\Omega)]^{2}$ and $\chi=(\chi_{1},\chi_{2})\in[L^{2}(\Omega)^{N}]^{2},$ we consider the optimization problem
$\mathcal{N}_{\pi,g}^{\chi}[\mu]:=\inf_{\rho,\phi}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]-[(\sigma\>\phi+\chi)\cdot\nu,g]_{\Gamma_{D}}\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi}[\mu-\rho]\right\},$
where
$\mathcal{A}_{\pi}^{\chi}[\mu-\rho]:=\mathcal{A}_{\pi}^{\chi}[\mu-\rho]_{1}\times\mathcal{A}_{\pi}^{\chi}[\mu-\rho]_{2}.$
See that $\mathcal{N}_{\pi,g}^{\chi}[\mu]$ may be written as
$\mathcal{N}_{\pi,g}^{\chi}[\mu]=\inf_{\rho,\omega}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\underbrace{\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]_{1}+\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]_{2}}_{\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]}\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2}\right\},$
where
(3.33) $\mathcal{I}_{\pi,g}^{\chi}[\mu]_{k}:=\inf_{\omega}\left\{\frac{1}{2}\int_{\Omega}\sigma_{k}\>|\omega|^{2}-\langle(\sigma\>\omega+\chi_{k})\cdot\nu,g_{k}\rangle_{\Gamma_{D_{k}}}\>:\>\omega\in\mathcal{A}_{\pi}^{\chi}[\mu]_{k}\right\}.$
Here the quantity $\mathcal{I}_{\pi,g}^{\chi}(\mu-\rho)$ represents the transition work associated with the distribution of mass $\mu-\rho$, and the term $\nabla\cdot\chi$ can be utilized to represent a significant external influence exerted on the system by an external force, $\chi$ (along with boundary conditions).
Thanks to [19], remember that
Proposition 2
(cf. [19]) For any $k=1,2,$ we have
(3.34) $\mathcal{I}_{\pi,g}^{\chi}[\mu]_{k}=\max_{z\in H^{1}(\Omega),\>z_{/\Gamma_{D_{k}}}=g_{k}}\left\{\int_{\Omega}\mu_{k}\>z-\int_{\Omega}\chi_{k}\cdot\nabla z-\ \int_{\Omega}F_{\sigma}[\nabla z]+\langle\pi_{k},z\rangle_{\Gamma_{N_{k}}}\right\}.$
Moreover, $\omega$ and $z$ are optimal if and only if $\omega=\nabla z$ and $z$ is the unique solution of the following PDE
(3.35) $\left\{\begin{array}[]{ll}-\nabla\cdot(\sigma_{k}\>\nabla z)=\mu_{k}-\nabla\cdot\chi_{k}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla z+\chi_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ z=g_{k},&\hbox{ on }\Gamma_{D_{k}}.\end{array}\right.$
Coming back to the problem $\mathcal{N}_{\pi,g}^{\chi}[\mu],$ we have the following duality result.
Theorem 3
For any $\mu=(\mu_{1},\mu_{2})\in[L^{2}(\Omega)]^{2}$ and $\chi=(\chi_{1},\chi_{2})\in[L^{2}(\Omega)^{N}]^{2},$ we have :
(3.36) $\mathcal{N}_{\pi,g}^{\chi}[\mu]=\underbrace{\max_{\eta\in[H^{1}(\Omega)]^{2},\eta_{/\Gamma_{D}}=g}\left\{\int_{\Omega}\mu\cdot\eta+\int_{\Omega}[\chi\cdot\nabla\eta]-\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}\beta_{\alpha}^{*}[\eta]+[\pi,\eta]_{\Gamma_{N}}\right\}}_{=:D_{\pi,g}^{\chi}[\mu]},$
where
$\beta_{\alpha}^{*}[\eta]:=\beta^{*}\left(x,\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\right),\quad\hbox{ for any }\eta\in[L^{2}(\Omega)]^{2}.$
Moreover, $(\rho,\phi)$ and $\eta$ are solutions of $\mathcal{N}_{\pi,g}^{\chi}[\mu]$ and $D_{\pi,g}^{\chi}[\mu],$ respectively, if and only if for each $k=1,2,$ $\phi_{k}=\nabla\eta_{k}$ and the couple $(\rho,\eta)$ satisfies the following system of PDE :
(3.37) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=\mu_{k}+\nabla\cdot\chi_{k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k}&\hbox{ on }\Gamma_{D_{k}},\end{array}\right.$
in the sense that $\rho_{k}\in L^{2}_{+}(\Omega),$ $\eta_{k}\in H^{1}(\Omega),$ $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$ for $k=1,2,$ and we have
  • $\rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0,$ a.e. in $\Omega,$
  • $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})$, a.e. in $\Omega,$
  • For each $k=1,2,$
    $\int_{\Omega}\rho_{k}\>\xi_{k}+\int_{\Omega}(\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nabla\xi_{k}=\int_{\Omega}\mu_{k}\>\xi_{k}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{k}}.$
    for any $\xi_{k}\in H^{1}_{\Gamma_{{}_{k}}}(\Omega).$
To the proof of this theorem, we see first that we have.
Proposition 3
For any $\mu=(\mu_{1},\mu_{2})\in[L^{2}(\Omega)]^{2}$ and $\chi\in[L^{2}(\Omega)^{N}]^{2},$ the problems $\mathcal{N}_{\pi,g}^{\chi}[\mu]$ and $D_{\pi,g}^{\chi}[\mu]$ have solutions.
Proof :
Let $(\rho^{n},\phi^{n})$ be a minimizing sequence $\mathcal{N}_{\pi,g}^{\chi}[\mu].$ Thanks to the assumptions on $\beta$, it is clear that $\rho^{n}$ and $\phi^{n}$ are bounded in $[L^{2}_{+}(\Omega)]^{2}$ and $[L^{2}(\Omega)^{N}]^{2}$, respectively. So, there exists a subsequence that we denote again by $(\rho^{n},\phi^{n})\in[L^{2}_{+}(\Omega)]^{2}\times[L^{2}(\Omega)^{N}]^{2}$, and $(\rho,\phi)\in[L^{2}_{+}(\Omega)]^{2}\times[L^{2}(\Omega)^{N}]^{2},$ such that
$\rho^{n}\to\rho,\quad\hbox{ in }[L^{2}(\Omega)]^{2}\hbox{-weak}^{*}$
and
[…]
Combining, in addition, (2.12) with the fact that $-[\nabla\cdot(\sigma\>\phi^{n}+\chi)]=\mu-\rho^{n}$ and $(\sigma\>\phi^{n}+\chi)\cdot\nu=\pi$ on $\Gamma_{N},$ we see that
$[(\sigma\>\phi^{n}+\chi)\cdot\nu,g]_{\Gamma_{D}}\to[(\sigma\>\phi+\chi)\cdot\nu,g]_{\Gamma_{D}}.$
Clearly $(\rho,\phi)$ is an admissible test function for the optimization problem $\mathcal{N}_{\pi,g}^{\chi}[\mu]$ ; i.e. $(\rho,\phi)\in\mathcal{A}_{\pi}^{\chi}[\mu-\rho]$. Then, using the lsc and convexity of $\beta$ and $F_{\sigma}$, we deduce that $(\rho,\phi)$ is a solution of the optimization problem $\mathcal{N}_{\pi,g}^{\chi}[\mu].$ The proof for $D_{\pi,g}^{\chi}[\mu]$ follows more or less the same ideas. $\square$
To prove Theorem 3, let us consider the application $K\>:\>[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}\to I\!\!R$ given by
$K[f]:=\int_{\Omega}f_{0}\cdot\tilde{g}-\int_{\Omega}[\overline{f}\cdot\nabla\tilde{g}]+\inf_{\rho,\phi}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]\>dx-[(\sigma\>\phi+\chi+\overline{f})\cdot\nu,g]_{\Gamma_{D}}\>\right.$
$\left.:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}(\mu_{k}+f_{0}-\rho)\right\},$
where $f_{0}=(f_{01},f_{02})$ and $\overline{f}=(\overline{f}_{1},\overline{f}_{2})$ are given by the decomposition $f_{k}=f_{0k}+\nabla\cdot\overline{f}_{k}$ in $H^{-1}_{\Gamma_{D_{k}}}(\Omega),$ for each $k=1,2.$ Thank to (2.12), we see that
$[(\sigma\>\phi+\chi+\overline{f})\cdot\nu,g]_{\Gamma_{D}}=\int_{\Omega}[(\sigma\>\phi+\chi+\overline{f})\cdot\nabla\tilde{g}]-\int_{\Omega}(f_{0}+\mu-\rho)\cdot\tilde{g},$
for any $\rho\in[L^{2}_{+}(\Omega)]^{2}$ and $\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}[\mu+f_{0}-\>\rho].$ This implies that $K$ may be written as
(3.38) $\begin{array}[]{c}K[f]=\inf_{\rho,\phi}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]\>dx-\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]+\int_{\Omega}(\mu-\rho)\cdot\tilde{g}\>\right.\\ \\ \>\left.:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}[\mu+f_{0}-\rho]\right\}.\end{array}$
The following duality result is the key to proving Theorem 3. After, we need simply to compute the Legendre transform of $K$ to complete the proofs.
Lemma 1
We have
(3.39) $-K[0]=-K^{**}[0]=\min_{\eta\in W^{1,p}(\Omega)}K^{*}[\eta],$
Proof :
To prove (3.39), it is enough to prove that $K$ is convex and l.s.c. and conclude by classical duality results (cf. [14]).
Convexity : For any $f,\>h\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ taking $(\rho_{f},\phi_{f})$ and $(\rho_{h},\phi_{h})$ the solutions corresponding to the optimization problems $K[f]$ and $K[h]$ respectively, one sees that
$\underbrace{t\phi_{h}+(1-t)\phi_{f}}_{\phi_{t}}\in\mathcal{A}_{\pi}^{\overline{h}_{t}}[\mu-\alpha\underbrace{(t\rho_{h}+(1-t)\rho_{f}}_{\rho_{t}}+\underbrace{(th_{0}+(1-t)f_{0})}_{h_{0t}}],$
where $\overline{h}_{t}=t\overline{h}+(1-t)\overline{f},$ for any $t\in[0,1].$ Moreover, using the convexity of $\beta(x,.)$ and $F_{\sigma},$ we have
$\int\beta_{\alpha}[\rho_{t}]+\>\int_{\Omega}F_{\sigma}[\phi_{t}]-[(\sigma\>\phi_{t}+\chi+\overline{f}_{t})\cdot\nu,g]_{\Gamma_{D}}\leq t\left(\int_{\Omega}\beta_{\alpha}[\rho_{h}]+\>\int_{\Omega}F_{\sigma}[\phi_{h}]-[\phi_{h}+\chi+\overline{h})\cdot\nu,g]_{\Gamma_{D}}\right)$
$+(1-t)\left(\int_{\Omega}\beta_{\alpha}[\rho_{f}]+\>\int_{\Omega}F_{\sigma}[\phi_{f}]-[(\sigma\>\phi_{f}+\chi+\overline{f})\cdot\nu,g]_{\Gamma_{D}}\right),$
which implies that
$K[th+(1-t)f]\leq tK[h]+(1-t)K[f].$
Lower semi-continuity : Let us consider $f^{n}$ a sequence of $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}$ which converges to $f.$ That is a sequence of $[L^{2}(\Omega)]$ and $[L^{2}(\Omega)^{N}]$ functions $f_{0}^{n}$ and $\overline{f}^{n}$ respectively, such that
$[f^{n},\xi]_{\Omega}=\int_{\Omega}f_{0}^{n}\cdot\xi-\int_{\Omega}[\overline{f}^{n}\cdot\nabla\xi]\to\int_{\Omega}f_{0}\cdot\xi-\int_{\Omega}[\overline{f}\cdot\nabla\xi]=[f,\xi]_{\Omega},\quad\forall\xi\in[H^{1}_{\Gamma_{D}}(\Omega)]^{2}.$
Let us prove that $K[f]\leq\liminf_{n\to\infty}K[f^{n}].$ To this aim, we consider $(\rho^{n},\phi^{n})$ be the solution corresponding to $K[f^{n}]$ ; i.e.
(3.40) $\phi^{n}\in\mathcal{A}_{\pi}^{\chi+\overline{f}^{n}}[\mu+f_{0}^{n}-\>\rho^{n}].$
and
[…]
where we use (3.38). We can assume that $K[f^{n}]$ is bounded. So, there exists $C<\infty$ such that
[…]
Using assumptions $(H1)$ and $(H2)$ with Young formula, we see that $\rho^{n}$ and $\phi^{n}$ are bounded in $[L^{2}_{+}(\Omega)]^{2}$ and $[L^{2}(\Omega)^{N}]^{2}$, respectively. So, there exists $(\rho,\phi)\in[L^{2}_{+}(\Omega)]^{2}\times[L^{2}(\Omega)^{N}]^{2}$ and a sub-sequence that we denote again by $(\rho^{n},\phi^{n})$, such that
[…]
and
[…]
Using (3.40), we have
(3.42) $\int_{\Omega}[(\sigma\>\phi^{n}-\chi)\cdot\nabla\xi]=\int_{\Omega}(\mu-\rho^{n})\cdot\xi+[f^{n},\xi]_{\Omega}+[\pi,\xi]_{\Gamma_{N}},\quad\forall\xi\in[H^{1}_{\Gamma_{D}}(\Omega)]^{2},$
and, by letting $n\to\infty,$ we get
(3.43) $\int_{\Omega}[(\sigma\>\phi-\chi)\cdot\nabla\xi]=\int_{\Omega}(\mu-\rho)\cdot\xi+[f,\xi]_{\Omega}+[\pi,\xi]_{\Gamma_{N}},\quad\forall\xi\in[H^{1}_{\Gamma_{D}}(\Omega)]^{2}.$
This implies that $\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}[\mu+f_{0}-\>\rho].$ Then, using the l.s.c. and convexity of $\beta$ and $F,$ we have
$\displaystyle K[f]$$\leq$ $\displaystyle\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]-\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]+\int_{\Omega}(\mu-\rho)\cdot\tilde{g}$
[…]
$\displaystyle$== $\displaystyle\liminf_{n\to\infty}K[f^{n}].$
Thus the result. $\square$
Proof of Theorem 3
: We only need to compute $K_{1}^{*}.$ By definition, for any $\eta\in[H^{1}_{\Gamma_{D}}(\Omega)]^{2},$ we have
$K^{*}[\eta]=\max\left\{[f,\eta]_{\Omega}-K[f]\>:\>f\in[H^{1}_{\Gamma_{D}}(\Omega)]^{2}\right\}$
$=\max_{f_{0},\overline{f},\phi}\Big{\{}\underbrace{\int_{\Omega}f_{0}\cdot\eta-\int_{\Omega}[\overline{f}\cdot\nabla\eta]-\int_{\Omega}\beta_{\alpha}[\rho]-\>\int_{\Omega}F_{\sigma}[\phi]+\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]-\int_{\Omega}(\mu-\rho)\cdot\tilde{g}}_{H[f_{0},\overline{f},\rho,\phi]}$
$\>:\>(f_{0},\overline{f})\in[L^{2}(\Omega)]^{2}\times[L^{2}(\Omega)^{N}]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}[\mu-\rho+f_{0}],\>\rho\in[L^{2}_{+}(\Omega)]^{2}\Big{\}}.$
Using (2.12), we have
$\int_{\Omega}f_{0}\cdot\eta-\int_{\Omega}[\overline{f}\cdot\nabla\eta]=\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\eta]-\int_{\Omega}(\mu-\rho)\cdot\eta-[\pi,\eta]_{\Gamma_{N}},$
so that
$\displaystyle H[f_{0},\overline{f},\rho,\phi]$ $\displaystyle$== $\displaystyle\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta_{\alpha}[\rho])$
$\displaystyle-\int_{\Omega}\mu\cdot(\eta+\tilde{g})+\int_{\Omega}[\chi\cdot\nabla(\eta+\tilde{g})]-[\pi,\eta]_{\Gamma_{N}}.$
This implies that
$\displaystyle K^{*}[\eta]$ $\displaystyle$== $\displaystyle\max_{\rho,\phi}\left\{\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta_{\alpha}[\rho])-\int_{\Omega}\mu\cdot(\eta+\tilde{g})]\right.$
$\displaystyle\hskip 85.35826pt\left.+\int_{\Omega}\chi\cdot\nabla(\eta+\tilde{g})-[\pi,\eta]_{\Gamma_{N}}\>:\>(\rho,\phi)\in[L^{2}_{+}(\Omega)]^{2}\times[L^{2}(\Omega)^{N}]^{2}\right\}$
$\displaystyle$== $\displaystyle\max_{\phi\in[L^{2}(\Omega)^{N}]^{2}}\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\max_{\rho\in[L^{2}_{+}(\Omega)]^{2}}\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta_{\alpha}[\rho])$
$\displaystyle\left.\hskip 85.35826pt-\int_{\Omega}\mu\cdot(\eta+\tilde{g})+\int_{\Omega}[\chi\cdot\nabla(\eta+\tilde{g})]-[\pi,\eta]_{\Gamma_{N}}\right\}.$
Using the assumptions $(H1)$ and $(H2)$, we deduce that
$\displaystyle K^{*}[\eta]$ $\displaystyle$== $\displaystyle\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}\beta_{\alpha}^{*}[\eta+\tilde{g}]-\int_{\Omega}\mu\cdot(\eta+\tilde{g})+\int_{\Omega}[\chi\cdot\nabla(\eta+\tilde{g})]-[\pi,\eta]_{\Gamma_{N}}.$
Combining this with (3.39), we obtain
$\displaystyle-K[0]$ $\displaystyle$== $\displaystyle\min_{\eta\in[H^{1}(\Omega)]^{2}}\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}\beta_{\alpha}^{*}[\eta+\tilde{g}]-\int_{\Omega}\mu\cdot(\eta+\tilde{g})+\int_{\Omega}[\chi\cdot\nabla(\eta+\tilde{g})]-[\pi,\eta]_{\Gamma_{N}}$
$\displaystyle$== $\displaystyle\min_{\eta\in[H^{1}(\Omega)]^{2},\>\eta_{/\Gamma_{D}}=g}\int_{\Omega}F_{\sigma}[\nabla\eta]+\int_{\Omega}\beta_{\alpha}^{*}[\eta]-\int_{\Omega}\mu\cdot\eta+\int[\chi\cdot\nabla\eta]-[\pi,\eta]_{\Gamma_{N}}.$
Thus the duality $\mathcal{N}_{\pi,g}^{\chi}(\mu)=D_{\pi,g}^{\chi}(\mu).$ Now, thanks to Lemma 3, let us consider $(\rho,\phi)$ and $\eta$ be the solutions of $\mathcal{N}_{\pi,g}^{\chi}[\mu]=D_{\pi,g}^{\chi}[\mu].$ We have
$\int_{\Omega}\mu\cdot\eta-\int_{\Omega}[\chi\cdot\nabla\eta]-\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}\beta_{\alpha}^{*}[\eta]+[\pi,\eta]_{\Gamma_{N}}$
$=\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]-[(\sigma\>\phi+\chi)\cdot\nu,g]_{\Gamma_{D}}.$
Since $\phi\in\mathcal{A}_{\pi}^{\chi}[\mu-\rho],$ $\eta\in[H^{1}(\Omega)]^{2}$ and $\eta=g$ on $\Gamma_{D},$ we also have
$\int_{\Omega}\rho\cdot\eta+\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\eta]=\int_{\Omega}\mu\cdot\eta+[\pi,\eta]_{\Gamma_{N}}+[(\sigma\>\phi+\chi)\cdot\nu,g]_{\Gamma_{D}}.$
Combining both equation, we get
$\int_{\Omega}F_{\sigma}[\phi]+\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}[\sigma\>\phi\cdot\nabla\eta]=\int_{\Omega}\eta\cdot\rho-\int_{\Omega}\beta_{\alpha}^{*}[\eta]-\int_{\Omega}\beta[\rho].$
This implies that, $\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}F_{\sigma}[\phi]=[\phi\cdot\nabla\eta]$ and $\beta_{\alpha}^{*}[\eta]+\beta[\rho]=\rho\cdot\eta$ , a.e. in $\Omega,$ for each $k=1,2.$ Thus $(\rho,\eta,\phi)$ is a solution of the PDE (3.44). For the the proof of the converse part, one sees first directly that $D_{\pi,g}^{\chi}[\mu]\leq\mathcal{N}_{\pi,g}^{\chi}[\mu].$ Then, by working with the solution of (3.44) one proves $D_{\pi,g}^{\chi}[\mu]=\mathcal{N}_{\pi,g}^{\chi}[\mu].$
$\square$
To end up this section, we prove the following results which will be useful for the sequel
Corollary 1
For any $\mu=(\mu_{1},\mu_{2})\in[L^{2}(\Omega)]^{2}$, $\pi=(\pi_{1},\pi_{2})\in[H^{-1/2}(\Gamma_{N})]$ and $g=(g_{1},g_{2})\in[H^{1/2}_{00}(\Gamma_{D})]$ the system of PDE
(3.44) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=\mu_{k}-\nabla\cdot\chi_{k}\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{ for }k=1,2,$
has a solution $(\rho,\eta)$ in the sense that $\rho_{k}\in L^{2}_{+}(\Omega),$ $\eta_{k}\in H^{1}(\Omega),$ $\eta_{k}=g_{1}$ on $\Gamma_{D_{k}},$ for $k=1,2,$ $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),$ $\rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0$ a.e. in $\Omega,$ and for each $k=1,2,$
$\int_{\Omega}\rho_{k}\>\xi_{k}+\int_{\Omega}\sigma_{k}\>\nabla\eta_{k}\cdot\nabla\xi_{k}=\int_{\Omega}\mu_{k}\>\xi_{k}-\int\chi_{k}\cdot\nabla\xi_{k}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{N_{k}}},\quad\forall\>\xi\in H^{1}_{\Gamma_{D_{1}}}(\Omega)\times H^{1}_{\Gamma_{D_{2}}}(\Omega).$
In particular, we have
(3.45) $\int_{\Omega}\rho\cdot(\eta-\tilde{g})+\int_{\Omega}[\sigma\>\nabla\eta\cdot\nabla(\eta-\tilde{g})]=\int_{\Omega}\mu\cdot(\eta-\tilde{g})-\int[\chi\cdot\nabla(\eta-\tilde{g})]+[\pi,(\eta-\tilde{g})]_{\Gamma_{N}}.$
Remark 5
  1. 1.
    For simplicity, we present in this paper the case where $F_{\sigma}$ is defined as in (3.31). However, the framework can be extended to a more general form: $F[\Phi]=F_{1}(x,\phi_{1})+F_{2}(x,\phi_{2}),$ where each each $F_{k}$ is tailored to capture the specific diffusion properties of the specie $k=1,2.$ Readers interested in exploring the expected associated dynamic for general case for $F$ can refer to [19].
  2. 2.
    While assumption (2.7) suffices for the proofs in this paper, exploring specific scenarios where either $\sigma_{1}\equiv 0$ or $\sigma_{2}\equiv 0$ holds significant interest for applications. We believe that some of the results might still be valid under the assumptions $\min_{x\in\overline{\Omega}}\sigma_{1}(x)\>\min_{x\in\overline{\Omega}}\sigma_{2}(x)\neq 0$, potentially requiring additional compatibility conditions. We defer the details of this case for future work focused on applications.
4 Existence for evolution problem
Our objective here is to establish the existence of a weak solution. As mentioned in the introduction, we proceed by discretizing time with discrete steps of size $\tau$, denoted as $t_{0}=0<t_{1}<\cdots<t_{n}<T$, and then consider the limit as $\tau$ approaches zero in the sequence of piecewise constant curves
$\rho^{\tau}=\sum_{i}\rho^{i}\>\chi_{[t_{k},t_{i+1}[}.$
Here $\rho^{i}$ is given, for each $i=1,....n,$ by
$\rho^{i}=\hbox{argmin}_{\rho}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\tau\>\int_{\Omega}F_{\sigma}[\phi]\>dx-\tau\>[(\sigma\>\phi+\rho^{i-1}V^{i}+\overline{f}^{i})\cdot\nu,g]_{\Gamma_{D}}\right.$
$\left.\>:\>(\rho,\phi)\in\mathcal{A}_{\pi}^{\rho^{i-1}V^{i}+\overline{f}^{i}}[(\tau f_{0}^{i}+\rho^{i-1}-\rho)/\tau],\>\rho\in[L^{2}_{+}(\Omega)]^{2}\right\}$
and
$f_{0}^{i}(.):=\left(\frac{1}{\tau}\int_{t_{i-1}}^{t_{i}}f_{01}(t,.)\>dt,\frac{1}{\tau}\int_{t_{i-1}}^{t_{i}}f_{02}(t,.)\>dt\right)\hbox{ and }\overline{f}^{i}(.):=\left(\frac{1}{\tau}\int_{t_{i-1}}^{t_{i}}\overline{f}_{1}(t,.)\>dt,\frac{1}{\tau}\int_{t_{i-1}}^{t_{i}}\overline{f}_{2}(t,.)\>dt\right),$
In other words, $\rho^{i}$ is given by
$\rho^{i}=\hbox{argmin}_{\rho}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\tau\>\mathcal{I}_{\pi,g}^{\rho^{i-1}V^{i}+\overline{f}^{i}}[(\tau f_{0}^{i}+\rho^{i-1}-\rho)/\tau]\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2}\right\}.$
Thanks to Theorem 3, we know that $\rho_{i}$ satisfies the following system PDE
(4.46) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}^{i}-\tau\>\nabla\cdot(\nabla\eta_{k}^{i}+\rho^{i-1}_{k}V_{k}^{i}+\overline{f}_{k}^{i})=\tau f_{0k}^{i}+\rho_{k}^{i-1},\\ \\ \rho^{i}_{1},\>\rho^{i}_{2}\geq 0,\>\frac{\eta^{i}_{1}}{\alpha_{1}}\vee\frac{\eta^{i}_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho^{i}_{1}+\alpha_{2}\rho^{i}_{2})\\ \\ \rho^{i}_{1}\left(\frac{\eta^{i}_{2}}{\alpha_{2}}-\frac{\eta^{i}_{1}}{\alpha^{i}_{1}}\right)^{+}=\rho^{i}_{2}\left(\frac{\eta^{i}_{1}}{\alpha_{1}}-\frac{\eta^{i}_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}^{i}+\rho^{i-1}_{k}V_{k}^{i}+\overline{f}_{k}^{i})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}^{i}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{ for }k=1,2,$
in the sense that $\rho_{k}^{i}\in L^{2}_{+}(\Omega),$ $\eta_{k}^{i}\in H^{1}(\Omega),$ $\eta_{k}^{i}=g_{k}$ on $\Gamma_{D_{k}},$ for $k=1,2$ and we have
  • $\frac{\eta^{i}_{1}}{\alpha_{1}}\vee\frac{\eta^{i}_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho^{i}_{1}+\alpha_{2}\rho^{i}_{2})$ $\rho^{i}_{1}\left(\frac{\eta^{i}_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha^{i}_{1}}\right)^{+}=\rho^{i}_{2}\left(\frac{\eta^{i}_{1}}{\alpha_{1}}-\frac{\eta^{i}_{2}}{\alpha_{2}}\right)^{+}=0$ a.e. in $\Omega.$
  • For each $k=1,2,$
[…]
    for any $\xi\in H^{1}_{\Gamma_{D_{1}}}(\Omega)\times H^{1}_{\Gamma_{D_{2}}}(\Omega).$
Let us define the sequence
$\tilde{\rho}^{\tau}(t)=\frac{(t-t_{i})\rho^{i+1}-(t-t_{i+1})\rho^{i}}{\tau},\quad\hbox{ for any }t\in[t_{i},t_{i}+1),\quad i=0,....n-1.$
In particular one sees that
(4.47) $\tilde{\rho}^{k}(t)-\rho^{\tau}(t)=(t-t_{i})\partial_{t}\rho^{\tau}(t),\quad\hbox{ for any }t\in[t_{i},t_{i+1}),\>i=0,1,...n-1,$
and for a.e. $t\in(0,T),$ the triplet $(\tilde{\rho}^{\tau},\rho^{\tau},p^{\tau})$ satisfies the following PDE (in the sense of Corollary 1)
(4.48) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\partial_{t}\tilde{\rho}_{k}^{\tau}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{i-1}_{k}V_{k}^{\tau}+\overline{f}_{k}^{\tau})=f_{0k}^{\tau},\\ \\ \rho^{\tau}_{1},\>\rho^{\tau}_{2}\geq 0,\>\frac{\eta^{\tau}_{1}}{\alpha_{1}}\vee\frac{\eta^{\tau}_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho^{\tau}_{1}+\alpha_{2}\rho^{\tau}_{2})\\ \\ \rho^{i}_{1}\left(\frac{\eta^{\tau}_{2}}{\alpha_{2}}-\frac{\eta^{\tau}_{1}}{\alpha^{\tau}_{1}}\right)^{+}=\rho^{i}_{2}\left(\frac{\eta^{\tau}_{1}}{\alpha_{1}}-\frac{\eta^{\tau}_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{i-1}_{k}V_{k}^{\tau}+\overline{f}_{k}^{\tau})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}^{\tau}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{for }k=1,2,$
The next step is to analyze the behavior of the approximation triplet $(\rho^{\tau},\tilde{\rho}^{\tau},\eta^{\tau})$ as the parameter $\tau\to 0.$ We aim to proof that, selecting a subsequence if necessary, the limit of this triplet converges to a solution of the system of PDEs given by equation (2.13). This objective motivates the following proposition, whose proof is deferred to the end of this section.
Proposition 4
There exists $\rho\in L^{\infty}\left(0,T;[L^{2}(\Omega)]^{2}\right)$, $\eta\in L^{2}\left(0,T;[H^{1}(\Omega)]^{2}\right),$ such that $\eta=g$ on $\Gamma_{D},$ and there exists subsequences that we denote by again $\rho^{\tau}$, $\tilde{\rho}^{\tau}$ and $p^{\tau},$ such that
[…]
Moreover, $(\rho,\eta)$ satisfies the state equations (2.14).
Then, the proof of Theorem 1 follows simply by passing to the limit in the weak formulation of (4.48).
Proof of Theorem 1
: Thanks to Proposition 4, it remains to proof that the couple $(\rho,\eta)$ satisfies the weak formulation (2.15). Remember that, for any $t\in(0,T)$ and $\xi\in[H^{1}_{\Gamma_{D}}(\Omega)]^{2},$ we have
$\int_{\Omega}\partial_{t}\tilde{\rho}^{\tau}(t)\cdot\xi+\int_{\Omega}[\sigma\>\nabla\eta^{\tau}(t)\cdot\nabla\xi]=[f^{\tau}(t),\xi]_{\Omega}-\int_{\Omega}[(\rho^{\tau}(t-\tau)V^{\tau}(t)\cdot\nabla\xi]+[\pi,\xi]_{\Gamma_{N}},$
and then
(4.52) $\begin{array}[]{c}\frac{d}{dt}\int_{\Omega}\tilde{\rho}^{\tau}(t)\cdot\xi+\int_{\Omega}[(\sigma\>\nabla\eta^{\tau}(t)+\rho^{\tau}(t)V^{\tau}(t))\cdot\nabla\xi]=[f^{\tau}(t),\xi]_{\Omega}+\langle\pi,\xi\rangle_{\Gamma_{N}}\\ \\ +\underbrace{\int_{\Omega}[(\rho^{\tau}(t)-\rho^{\tau}(t-\tau))V^{\tau}(t)\cdot\nabla\xi]}_{J_{\tau}(t)},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T)).\end{array}$
Using the fact that $\rho^{\tau}$ is bounded in $[L^{2}(Q)]^{2}$ and $V^{\tau}$ is relatively compact in $[L^{2}(Q)^{N}]^{2},$ we see that $J_{\tau}\to 0$ weakly. So, passing to the limit in (4.52), we deduce that the couple $(\rho,\eta)$ is a weak solution of the problem (2.13). $\square$
Now, let us prove Proposition 4. We begin by to prove the following.
Lemma 2
For any $t\in[0,T),$ we have
(4.53) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{\tau}(t)]-\rho^{\tau}(t)\cdot\tilde{g})+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\int_{0}^{t}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ \hskip 85.35826pt-\int_{0}^{t}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{0}^{t}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ for any }t\in[0,T).\end{array}$
Proof :
Thanks to (3.45), we know that
(4.54) $\begin{array}[]{c}\int_{\Omega}\rho^{i}\cdot(\eta^{i}-\tilde{g})+\tau\>\int_{\Omega}[\sigma\>\nabla\eta^{i}\cdot\nabla(\eta^{i}-\tilde{g})]=\int_{\Omega}\rho^{i-1}\cdot(\eta^{i}-\tilde{g})+\tau\>[f^{i},\eta^{i}-\tilde{g}]_{\Omega}\\ \\ -\tau\int_{\Omega}[\rho^{i-1}\>V^{i}\cdot\nabla(\eta^{i}-\tilde{g})]+\tau[\pi,\eta^{i}]_{\Gamma_{N}}.\end{array}$
Remember that $\frac{\eta_{1}^{i}}{\alpha_{1}}\vee\frac{\eta_{2}^{i}}{\alpha_{2}}\in\partial\beta(.,\alpha_{1}\rho_{1}^{i}+\alpha_{2}\rho_{2}^{i}).$ This implies that $(\rho^{i}-\rho^{i-1})\cdot\eta^{i}\geq\beta_{\alpha}[\rho^{i}]-\beta_{\alpha}[\rho^{i-1}],$ and then
(4.55) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{i}]-\rho^{i}\cdot\tilde{g})-\int_{\Omega}(\beta_{\alpha}[\rho^{i-1}]-\rho^{i-1}\cdot\tilde{g})+\tau\>\int_{\Omega}F_{\sigma}[\nabla(\eta^{i}-\tilde{g})]\leq\tau\>[f^{i},\eta^{i}-\tilde{g}]_{\Omega}\\ \\ -\tau\>\int_{\Omega}[\rho^{i-1}\>V^{i}\cdot\nabla(\eta^{i}-\tilde{g})]+\tau\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]+\tau\>[\pi,\eta^{i}]_{\Gamma_{N}}.\end{array}$
Adding this inequality for $i=1...l\>\leq n,$ we obtain
(4.56) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{\tau}(t_{l})]-\rho^{\tau}(t_{l})\cdot\tilde{g})+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ -\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T),\end{array}$
which implies
(4.57) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{\tau}(t_{l})]-\rho^{\tau}(t_{l})\cdot\tilde{g})+\int_{0}^{t_{l}}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\int_{0}^{t_{l}}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ -\int_{0}^{t_{l}}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{0}^{t_{l}}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T).\end{array}$
Since $l$ is arbitrary, we deduce the result of the lemma.
$\square$
Lemma 3
There exists a constant $C=C(p,N,\Omega)$ such that :
(4.58) $\begin{array}[]{c}\int_{\Omega}(\alpha_{1}\>\rho_{1}^{\tau}(t)+\alpha_{2}\>\rho_{2}^{\tau}(t)-M)^{+2}+\int_{0}^{T}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\\ \\ \leq\frac{B}{C}\left\{1+\frac{T\|V^{\tau}\|_{\infty}^{2}}{C}\>\exp\left(\frac{T\|V^{\tau}\|_{\infty}^{2}}{C}\right)\right\},\end{array}$
where $B$ is given by
$B=\|\tilde{g}\|_{[L^{2}(\Omega)]^{2}}^{2}+\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+T\>\int_{\Omega}(\rho_{01}+\rho_{02}-M)^{+2}$
$+\sup_{0<\tau<1}\|V^{\tau}\|_{\infty}^{2}+\int_{0}^{T}\!\!\|\pi\|_{[H^{-1/2}_{\Gamma_{N}}]^{2}}^{2}+\sup_{0<\tau<1}\int_{0}^{T}\!\!\left(\|f_{0}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}+\|\overline{f}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}\right)$
Proof :
Thanks to Lemma 2, for any $t\in[0,T),$ we have
(4.59) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{\tau}(t)]-\rho^{\tau}(t)\cdot\tilde{g})+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+T\>\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]\\ \\ +\underbrace{\int_{0}^{t}\!\!\int_{\Omega}|[\rho^{\tau}(.-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]|}_{J_{1}^{\tau}}+\underbrace{\int_{0}^{t}|[\pi,\eta^{\tau}]_{\Gamma_{N}}+[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}|}_{J_{2}^{\tau}}.\end{array}$
Using Young inequality with $\epsilon$ and $C_{\epsilon},$ we see that
$\displaystyle J_{1}^{\tau}$$\leq$ $\displaystyle\int_{0}^{t}\!\!\int_{\Omega}|\rho^{\tau}_{1}(.-\tau)V^{\tau}_{1}\cdot\nabla(\eta^{\tau}_{1}-\tilde{g}_{1})+\rho^{\tau}_{2}(.-\tau)V^{\tau}_{2}\cdot\nabla(\eta^{\tau}_{2}-\tilde{g}_{2})|$
[…]
Moreover, since $\frac{1}{2}|a|^{2}\leq(|a|-M)^{+2}+M^{2},$ we get
$\displaystyle J_{1}^{\tau}$$\leq$ $\displaystyle C_{\epsilon}\|V^{\tau}\|_{\infty}^{2}\int_{0}^{t}\!\!\int_{\Omega}(\alpha_{1}\rho_{1}(.-\tau)+\alpha_{2}\rho_{2}(.-\tau)-M)^{+2}+\epsilon\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]$
$\displaystyle+C_{\epsilon}M^{2}\>T\>\|V^{\tau}\|_{\infty}^{2}\>|\Omega|$
[…]
$\displaystyle+\tau\>\int_{\Omega}(\alpha_{1}\rho_{01}+\alpha_{2}\rho_{02}-M)^{+2}.$
It is not difficult to see also that, for any $\epsilon>0,$ we can find $C_{\epsilon}>0,$ such that
$J_{2}^{\tau}\leq C_{\epsilon}\>\int_{0}^{T}\!\!\left(\|\pi\|_{[H^{-1/2}_{\Gamma_{N}}]^{2}}^{2}+\|f_{0}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}+\|\overline{f}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}\right)+\epsilon\int_{0}^{t}\!\!\int_{\Omega}+C_{\epsilon}\>M^{2}\>T\>\|V^{\tau}\|_{\infty}^{2}|\Omega|[\nabla(\eta^{\tau}-\tilde{g})].$
Using moreover, the fact that
$\beta_{\alpha}[\rho^{\tau}]-\rho^{\tau}\cdot\tilde{g}\geq\frac{C}{2}(\alpha_{1}\rho^{\tau}_{1}+\alpha_{2}\rho^{\tau}_{2}-M)^{+2}-\frac{1}{2C}|\tilde{g}_{1}\vee\tilde{g}_{2}|^{2}\geq\frac{C}{2}(\alpha_{1}\rho^{\tau}_{1}+\alpha_{2}\rho^{\tau}_{2}-M)^{+2}-\frac{1}{2C}|\tilde{g}|^{2},$
(4.59) implies that
(4.60) $\begin{array}[]{c}\frac{C}{2}\int_{\Omega}(\alpha_{1}\rho^{\tau}_{1}(t)+\alpha_{2}\rho^{\tau}_{2}(t)-M)^{+2}+(1-2\epsilon)\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\frac{1}{2C}\|\tilde{g}\|_{[L^{2}(\Omega)]^{2}}^{2}+T\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]\\ \\ +\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\tau\>\int_{\Omega}(\alpha_{1}\rho_{01}+\alpha_{2}\rho_{02}-M)^{+2}+C_{\epsilon}\>\|V^{\tau}\|_{\infty}^{2}\int_{0}^{t}\!\!\int_{\Omega}(\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}-M)^{+2}\\ \\ +\int_{0}^{T}\!\!\left(\|\pi\|_{[H^{-1/2}_{\Gamma_{N}}]^{2}}^{2}+\|f_{0}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}+\|\overline{f}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}\right)+C_{\epsilon}\>M^{2}\>T\>\|V^{\tau}\|_{\infty}^{2}|\Omega|,\quad\hbox{for any }t\in[0,T).\end{array}$
Working with a fixed small $\epsilon,$ we can find a constant $C=C(p,N,T,\Omega)>0,$ such that
(4.61) $\begin{array}[]{c}\int_{\Omega}(\alpha_{1}\rho^{\tau}_{1}(t)+\alpha_{2}\rho^{\tau}_{2}(t)-M)^{+2}+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla\eta^{\tau}-\tilde{g}]\\ \\ \leq\frac{B}{C}+\frac{\|V^{\tau}\|_{\infty}^{2}}{C}\>\int_{0}^{t}\!\!\int_{\Omega}(\alpha_{1}\rho^{\tau}_{1}(t)+\alpha_{2}\rho^{\tau}_{2}(t)-M)^{+2}.\end{array}$
This implies that
(4.62) $\int_{\Omega}(\alpha_{1}\rho^{\tau}_{1}(t)+\alpha_{2}\rho^{\tau}_{2}(t)-M)^{+2}\leq\frac{B}{C}\>\exp\left(T\|V^{\tau}\|_{\infty}^{2}/C\right),$
and (4.58) follows by (4.61). $\square$
Proof of Proposition 4
: Thanks to (4.58) it is clear that $\rho^{\tau}$ and $\tilde{\rho}^{\tau}$ are bounded in $L^{\infty}\left(0,T;[L^{2}(\Omega)]^{2}\right),$ and $\eta^{\tau}$ is bounded in $L^{2}\left(0,T;[H^{1}(\Omega)]^{2}\right).$ This implies (4.49) and (4.51). On the other hand, since
$\partial_{t}\tilde{\rho}_{k}^{\tau}=f_{k0}^{\tau}+\nabla\cdot(\rho_{k}^{\tau}(.-\tau)V_{k}^{\tau}+\sigma_{k}\>\nabla\eta_{k}^{\tau}+\overline{f}_{k}^{\tau})\quad\hbox{ in }\Omega,$
with $(\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{\tau}(.-\tau)V_{k}^{\tau}+\overline{f}_{k}^{\tau})\cdot\nu=\pi_{k}$ on $\Sigma_{N_{k}},$ and by (4.58), $\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho_{k}^{\tau}(.-\tau)V_{k}^{\tau}+\overline{f}_{k}^{\tau}$ is bounded in $L^{2}(Q),$ we see that $\partial_{t}\tilde{\rho}_{k}^{\tau}$ is bounded in $L^{2}\left(0,T;H^{-1}_{\Gamma_{D_{k}}}(\Omega)\right).$ Combining this with (4.47) we deduce that $\tilde{\rho}^{\tau}$ and $\rho^{\tau}$ have the same $[L^{2}(Q)]-$weak limit. Thus (4.50). Let us prove now that the couple $(\rho,\eta)$ satisfies the state equations (2.14). Thanks again to (4.58), $\eta_{1}^{\tau}\vee\eta_{2}^{\tau}$ is bounded in $L^{2}\left(0,T;H^{1}(\Omega)\right),$ and there exists $\overline{\eta}\in L^{2}\left(0,T;[H^{1}(\Omega)]^{2}\right)$ such that, by taking a sub-sequence if necessary,
(4.63) $\frac{\eta_{1}^{\tau}}{\alpha_{1}}\vee\frac{\eta_{2}^{\tau}}{\alpha_{2}}\to\overline{\eta},\quad\hbox{ in }L^{2}\left(0,T;H^{1}(\Omega)\right)-\hbox{weak},$
which satisfies moreover
(4.64) $0\leq\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\leq\overline{\eta},\quad\hbox{ a.e. in }Q.$
So, using Weak Aubin’s type Lemma (cf. [24, 5]), we get
(4.65) $\int_{0}^{T}\!\!\int_{\Omega}\rho_{k}^{\tau}\>\eta_{k}^{\tau}\>\varphi\quad\xrightarrow[\tau\to 0]{}\quad\int_{0}^{T}\!\!\int_{\Omega}\rho_{k}\>\eta_{k}\>\varphi,\quad\forall\>\varphi\in\mathcal{D}(Q)\hbox{ and }k=1,2,$
and also
(4.66) $\int_{0}^{T}\!\!\int_{\Omega}(\alpha_{1}\rho_{1}^{\tau}+\alpha_{2}\rho_{2}^{\tau})\frac{\eta_{1}^{\tau}}{\alpha_{1}}\vee\frac{\eta_{2}^{\tau}}{\alpha_{2}}\>\varphi\quad\xrightarrow[\tau\to 0]{}\quad\int_{0}^{T}\!\!\int_{\Omega}(\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\>\overline{\eta}\>\varphi\quad\forall\>\varphi\in\mathcal{D}(Q).$
Combining this with the fact that $\rho_{1}^{\tau}\>\eta_{1}^{\tau}+\rho_{2}^{\tau}\>\eta_{2}^{\tau}=(\alpha_{1}\rho_{1}^{\tau}+\alpha_{2}\rho_{2}^{\tau})\frac{\eta_{1}^{\tau}}{\alpha_{1}}\vee\frac{\eta_{2}^{\tau}}{\alpha_{2}},$ we obtain
(4.67) $\rho_{1}\>\eta_{1}+\rho_{2}\>\eta_{2}=(\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\>\overline{\eta},\quad\hbox{ a.e. in }Q,$
and, by (4.64), we deduce that
(4.68) $\rho_{1}\>\eta_{1}+\rho_{2}\>\eta_{2}=(\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}},\quad\hbox{ a.e. in }Q.$
Moreover, since $\frac{\eta_{1}^{\tau}}{\alpha_{1}}\vee\frac{\eta_{2}^{\tau}}{\alpha_{2}}\in\partial\beta(.,\alpha_{1}\rho_{1}^{\tau}+\alpha_{2}\rho_{2}^{\tau}),$ a.e. in $Q,$ by using again weak Aubin’s type Lemma and (4.66), we see that
$\overline{\eta}\in\partial\beta(.,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),\quad\hbox{ a.e. in }Q.$
To finish the proof, let us justify that $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(.,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),$ a.e. in $Q.$ If $\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}\neq 0,$ (4.67) and (4.68) imply clearly that $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}=\overline{\eta}\in\partial\beta(.,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),$ a.e. in $Q.$ If $\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}=0,$ the result follows from the fact that $\partial\beta(.,0)\ni 0\leq\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}2\leq\overline{\eta}\in\partial\beta(.,0).$
$\square$
5 Connection with $H^{-1}$-like theory in the homogeneous case
The aim of this section is to prove Theorem 2. So, we assume that
$\pi=0\hbox{ on }\Gamma_{N},\quad g=0\hbox{ on }\Gamma_{D}\quad\hbox{ and }\quad V\equiv 0.$
Let us consider the functional $\mathcal{E}$ as defined in Section 2. Then, we consider the Hilbert space $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ endowed with the norm
$\|f\|_{H^{-1}_{\Gamma_{D}}(\Omega)}=\left(\|f_{1}\|_{H^{-1}_{\Gamma_{D_{1}}}}^{2}+\|f_{2}\|_{H^{-1}_{\Gamma_{D_{2}}}}^{2}\right)^{1/2},\quad\hbox{ for any }f=(f_{1},f_{2})\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$
and its associate inner product
$[f,g]_{\Omega}=\langle f_{1},g_{1}\rangle_{1,\Omega}+\langle f_{2},g_{2}\rangle_{k,\Omega},\quad\hbox{ for any }f,g\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}.$
Our main result in this section concerns the characterization of $\partial E$ in terms of cross-diffusion system like in (3.37).
Proposition 5
For any $(\rho,f)\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}\times[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}.$ Then, $f\in\partial E(\rho)$ if and only if $\rho\in[L^{2}_{+}(\Omega)]^{2}$ and $\rho$ solves the following system
(5.69) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=f_{0k}+\nabla\cdot\overline{f}_{k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{k})\cdot\nu=0&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k}&\hbox{ on }\Gamma_{D_{k}},\end{array}\right.$
in the sense that, there exists $\eta_{k}\in H^{1}(\Omega),$ such that $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$ for $k=1,2,$ and we have
  • $\rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0,$ a.e. in $\Omega,$
  • $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})$, a.e. in $\Omega,$
  • For each $k=1,2,$
    $\int_{\Omega}(\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nabla\xi_{k}=\int_{\Omega}\mu_{k}\>\xi_{k}-\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{k}}.$
    for any $\xi_{k}\in H^{1}_{\Gamma_{{}_{k}}}(\Omega).$
Proof :
Our aim is to prove that
(5.70) $\partial\mathcal{E}=\mathcal{A},$
where $\mathcal{A}$ is the operator define in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ by $f\in\mathcal{A}(\rho)$ if and only if $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ $\rho\in[L^{2}(\Omega)_{+}]^{2}$ and $\rho$ is a solution of (5.69), in the sense of Proposition 5. Thanks to Corollary 1, for any $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ there exists $\rho\in[L^{2}_{+}(\Omega)]^{2},$ such that $f=\rho+\mathcal{A}(\rho).$ On the other hand, since $\mathcal{E}$ is convex and l.s.c. in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ the operator $\partial E$ defines a maximal monotone graph in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}.$ So, for the proof of (5.70), it is enough to prove that $\mathcal{A}\subseteq\partial\mathcal{E}.$ To this aim, we consider $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}$ and $\rho\in[L^{2}(\Omega)_{+}]^{2}$ be such that $f\in\mathcal{A}(\rho),$ and we prove that
(5.71) $[f,z-\rho]_{\Omega}\leq\int_{\Omega}\beta_{\alpha}[z]-\int_{\Omega}\beta_{\alpha}[\rho]+,\quad\hbox{ for any }z\in[L^{2}(\Omega)_{+}]^{2}.$
Let us consider $\phi_{z}$ and $\phi_{\rho}$ given by
$\phi^{z}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}|\sigma_{k}\>\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega)=z_{k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega)\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\},$
and
$\phi^{\rho}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}|\sigma_{k}\>\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega)=\rho_{k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega)\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\}.$
Since $f_{k}$ satisfies (5.69), using the definition of .,.$\langle.,.\rangle_{k,\Omega}$, we have
$\displaystyle[f,z-\rho]_{\Omega}$ $\displaystyle$== $\displaystyle\langle f_{1},z_{1}-\rho_{1}\rangle_{\Omega}+\langle f_{2},z_{2}-\rho_{2}\rangle_{\Omega}$
$\displaystyle$== $\displaystyle\int_{\Omega}\sigma_{1}\>\nabla\eta_{1}\cdot(\phi^{z}_{1}+\phi^{\rho}_{1})+\int_{\Omega}\sigma_{2}\>\nabla\eta_{2}\cdot(\phi^{z}_{2}+\phi^{\rho}_{2})$
$\displaystyle$== $\displaystyle\int_{\Omega}(\eta_{1}\>z_{1}+\eta_{2}\>z_{2})-\int_{\Omega}(\eta_{1}\>\rho_{1}+\eta_{2}\>\rho_{2})$
Moreover, since $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2}),$ a.e. in $\Omega,$ we obtain
$\displaystyle[f,z-\rho]_{\Omega}$ $\displaystyle$== $\displaystyle\int_{\Omega}(\eta_{1}\>z_{1}+\eta_{2}\>z_{2})-\int_{\Omega}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}$
[…]
Thus (5.71).
$\square$
Proof of Theorem 2
: Since $\partial\mathcal{E}$ is a maximal monotone graph, the existence of a strong solution of (2.21) follows directly from the well-established theory of nonlinear semigroups in Hilbert spaces (see [9] for details). This allows to conclude that Theorem 2 holds true as a direct consequence of Proposition 5. $\square$